### v1

# Shadow in the Cache: Unveiling and Mitigating Privacy Risks of KV-cache in LLM Inference

Abstract
The Key-Value (KV) cache, which stores intermediate attention computations (Key and Value pairs) to avoid redundant calculations, is a fundamental mechanism for accelerating Large Language Model (LLM) inference. However, this efficiency optimization introduces significant yet underexplored privacy risks. This paper provides the first comprehensive analysis of these vulnerabilities, demonstrating that an attacker can reconstruct sensitive user inputs directly from the KV-cache. We design and implement three distinct attack vectors: a direct Inversion Attack, a more broadly applicable and potent Collision Attack, and a semantic-based Injection Attack. These methods demonstrate the practicality and severity of KV-cache privacy leakage issues. To mitigate this, we propose KV-Cloak, a novel, lightweight, and efficient defense mechanism. KV-Cloak uses a reversible matrix-based obfuscation scheme, combined with operator fusion, to secure the KV-cache. Our extensive experiments show that KV-Cloak effectively thwarts all proposed attacks, reducing reconstruction quality to random noise. Crucially, it achieves this robust security with virtually no degradation in model accuracy and minimal performance overhead, offering a practical solution for trustworthy LLM deployment.
I Introduction
Large Language Models (LLMs) have ignited a paradigm revolution in artificial intelligence [33, 20], profoundly impacting various domains and applications, such as machine translation [40], chatbots [37], code generation [11], and content creation [42]. However, the immense scale of these models, characterized by billions or even trillions of parameters, coupled with the need to process extensive input sequences and engage in multi-turn dialogues, presents a substantial challenge to their efficient deployment and inference. This computational demand often translates into high latency and resource consumption, hindering broader accessibility and real-time applicability [18].
To address the efficiency bottlenecks in LLM inference, researchers have proposed several optimization techniques [44, 35]. Among these, the Key-Value cache (KV-cache) mechanism has emerged as a crucial and widely adopted solution [24, 42]. During the autoregressive generation process typical of LLMs, the attention mechanism computes key (K) and value (V) matrices for each token based on its preceding tokens. The KV-cache stores these intermediate attention computations (the K and V pairs) for tokens that have already been processed within the input sequence. By reusing these cached K and V pairs for the generation of subsequent tokens, the KV-cache significantly reduces redundant computations. This dramatically accelerates inference speed and improves throughput, especially for tasks involving long contexts or interactive sessions, making LLMs more practical for real-world deployment.
Figure 1: LLM inference service workflow in a privacy-preserving setting and its associated security risks.
However, the storage and potential sharing of the KV-cache introduce significant yet underexplored privacy concerns [38, 36], as illustrated in Figure 1. This vulnerability stems from a critical trade-off made in production systems: while end-user communication with the LLM service is typically encrypted (represented by the black lines in our figure), the KV-cache itself is almost always processed, transmitted between compute nodes, and persisted in plaintext. This design choice is a concession to performance, as the unacceptable latency overhead from cryptographically securing the often gigabyte-scale KV-cache would violate the stringent demands of real-time inference. The feasibility of exploiting such plaintext data exposure is not merely theoretical. It was recently demonstrated by the LeftoverLocals attack [28]. This vulnerability in certain AMD GPUs enables an unprivileged attacker to read data remnants from the GPU’s shared on-chip memory. This provides a concrete method to intercept and exfiltrate plaintext KV-cache data originating from other users’ concurrent inference sessions, proving that sensitive conversational data is indeed at risk. Crucially, the KV-cache is inherently derived from and directly correlated with user inputs. Consequently, if an attacker gains access to the KV-cache, they could potentially reverse-engineer sensitive information in the user’s original prompts or private data [6, 17]. A critical research gap exists in understanding and mitigating the privacy risks associated with KV-caches in LLMs, necessitating further in-depth research.
In this paper, to bridge the research gap, we present the first comprehensive study on the privacy risks of KV-cache in LLM inference. Specifically, we primarily study and answer the following two research questions:
  • RQ1: Is an attacker able to reconstruct user inputs from the KV-caches?
To address this question, we conduct a systematic investigation, demonstrating that attacks against the KV-cache are not only feasible but also diverse and broadly applicable. We design and implement three distinct classes of targeted privacy-stealing attacks. Each attack exposes inherent privacy risks of KV-caches from a different perspective.
We first explore two relatively direct attack paths. The first is the KV-cache Inversion Attack, a method that attempts to leverage known model weight matrices to directly reverse-calculate the input from the KV-cache. The second, more general-purpose approach is the KV-cache Collision Attack. This attack reframes input reconstruction as a matching task based on forward computation: an attacker iteratively uses a local model instance to generate KV-caches for candidate inputs and compares them against the intercepted target KV-cache to find a match. Because this method does not rely on any reverse computation, it has broader applicability. Finally, we introduce a novel, semantic-based KV-cache Injection Attack. The core of this attack is to leverage the LLM’s powerful capability to understand and execute instructions. By appending a specific instruction, such as “Repeat the previous content,” to the end of an intercepted KV-cache context, an attacker can induce the model to “echo” or generalize the core information held within the KV-cache.
Collectively, these attacks reveal that privacy leakage from the KV-cache is not merely a theoretical concern. Their diversity and feasibility constitute a significant threat to real-world LLM-based applications, underscoring the urgent necessity of designing specialized and efficient privacy-preserving mechanisms for the KV-cache.
  • RQ2: Can defenders effectively and efficiently mitigate or prevent user privacy leakage from the KV-cache?
In response to this challenge, this paper provides an affirmative answer by first analyzing the shortcomings of existing privacy-preserving techniques. Conventional methods like full cryptographic encryption [2] or Homomorphic Encryption (HE) [22] introduce prohibitive computational overhead and latency, rendering them impractical for the high-throughput demands of LLM inference. Meanwhile, applying Differential Privacy (DP) [8] requires adding a level of noise that often severely degrades the model’s inference accuracy to an unacceptable degree. Even specialized solutions like KV-Shield [38], while lightweight, suffer from critical security flaws. Their fixed shuffling mechanism, as we analyzed in Section IV-A, is vulnerable to statistical analysis and is incompatible with modern LLM architectures that use features like Rotary Positional Embedding (RoPE). These limitations highlight the urgent need for a novel defense mechanism, .
Therefore, we propose KV-Cloak, a lightweight and secure mechanism for KV-cache obfuscation designed to overcome these challenges. At its core, KV-Cloak employs a reversible matrix-based obfuscation scheme that guarantees lossless model accuracy. Its security is multi-layered: it applies secret invertible linear transformations to obscure statistical properties, and critically, introduces a one-time random permutation matrix for each data block. This dynamic permutation prevents attackers from building stable algebraic relations across multiple queries. To further enhance performance, KV-Cloak utilizes operator fusion, an optimization where a portion of the secret obfuscation matrices is algebraically fused into the LLM’s attention layer weights offline. This shifts the primary computational cost away from the latency-sensitive online inference phase, striking an effective balance between robust security, lossless accuracy, and high performance.
Our contributions can be summarized as follows.
  • Revealing the privacy risks of KV-cache in LLM inference: We systematically investigate the privacy risks of the KV-cache by designing and implementing three distinct attacks, Inversion, Collision, and Injection Attacks.
  • Proposing a lightweight and effective method to mitigate privacy leakage: We propose KV-Cloak, a novel and practical defense mechanism that uses a lightweight, reversible matrix-based obfuscation scheme combined with operator fusion to protect the KV-cache without degrading model accuracy and with minimal performance overhead.
  • Conducting extensive experiments to evaluate attacks and defenses: We conduct a systematic evaluation to empirically demonstrate and quantify the feasibility of attacks that reconstruct user input from the KV-cache of state-of-the-art LLMs, establishing it as a practical and severe threat. We further prove that our proposed KV-Cloak is a highly practical solution that achieves robust security, near-lossless model fidelity, and high efficiency simultaneously. Our experiments show that KV-Cloak, with its negligible impact on accuracy and minimal latency overhead (mostly $10\%$), successfully resolves the stark trade-off between security and utility that plagues alternative approaches such as DP.
II Background and Related Work
II-A Transformer-based LLM Inference
Prevailing large language models, such as LLaMA [32], DeepSeek [20] and Qwen [37], are predominantly based on the Transformer decoder architecture [42].
Self-Attention Mechanism.
The core of the Transformer architecture is the self-attention mechanism. It captures dependencies between different positions within an input sequence, allowing the model to flexibly focus on various parts of the sequence to better understand the context. For an input sequence $X=(x_{1},\ldots,x_{n})$, the self-attention layer first applies linear transformations to $x_{i}$ at each position to obtain Query ($q$), Key ($k$), and Value ($v$) vectors,
$q_{i}=x_{i}W_{q}^{\top}{R}_{\Theta,i}^{d},\quad k_{i}=x_{i}W_{k}^{\top}{R}_{\Theta,i}^{d},\quad v_{i}=x_{i}W_{v}^{\top}.$ (1)
Here, $W_{q},W_{k},W_{v}$ are learnable weight matrices for the attention layer, and ${R}_{\Theta,i}^{d}$ denotes the Rotary Position Embedding (RoPE)[29]. Then, the self-attention layer computes the dot product of the current query $q_{i}$ with all preceding keys $k_{j}$. The results are scaled, and a softmax function is applied to obtain attention weights $a_{ij}$. These weights are then used to compute a weighted sum of the value vectors. Finally, this sum is passed through a linear projection layer $W_{o}$ to produce the output $o_{i}$ for the current position. This process is detailed in Eq. (2), where $d$ is the dimension of each attention head.
$a_{ij}=\frac{\exp\left(q_{i}{k_{j}}^{\top}/\sqrt{d}\right)}{\sum_{t=1}^{i}\exp\left(q_{i}{k_{t}}^{\top}/\sqrt{d}\right)},o_{i}=\left(\sum_{j=1}^{i}a_{ij}v_{j}\right)W_{o}^{\top}.$ (2)
This mechanism enables the model to dynamically weigh the importance of different parts of the input sequence, thereby capturing complex contextual dependencies.
Autoregressive LLM Inference. The essence of large language model inference is to model the probability distribution $P(x_{1},\ldots,x_{n})$ of a token sequence $x_{1},\ldots,x_{n}$. Due to the sequential nature of natural language, prevailing LLMs predominantly employ an autoregressive text generation approach in inference tasks. The joint probability of the entire sequence can typically be autoregressively decomposed [5] into a product of conditional probabilities,
$P(x_{1},\ldots,x_{n})=P(x_{1})P(x_{2}|x_{1})\cdots P(x_{n}|x_{1},\ldots,x_{n-1}).$ (3)
Specifically, the model generates the output sequence one token at a time. When generating each new token, the model considers not only the initial user-provided input but also all previously generated tokens as context. For example, when generating the token $x_{i}$, the model utilizes information from tokens $x_{1}$ to $x_{i-1}$. While this sequential generation method effectively captures dependencies within the sequence, it introduces significant computational redundancy. At each step of generating a new token, attention computations for all preceding tokens (particularly the calculation of $k$ and $v$) must be recalculated. This repetitive computation substantially impacts inference efficiency.
KV-cache in LLM Inference. In the context of autoregressive generation, the previously generated tokens form the context for predicting the next token. To avoid redundant computations, LLMs employ a caching strategy known as the KV-cache. For each processed token (either from the input or previously generated by the model), its corresponding $k$ and $v$ vectors, computed by the self-attention layers, are stored in the KV-cache. When generating the next token, the model reuses these cached $k$ and $v$ vectors from all preceding tokens, only needing to compute the $q,k,v$ vector for the current prediction step and the $k$ and $v$ vectors for the most recently generated token to add to the cache. This significantly accelerates the inference process, especially for long sequences.
II-B Privacy Attacks against LLMs
Inversion Attacks. While LLMs offer powerful capabilities, their inference process exposes new attack surfaces that can lead to the theft of sensitive user inputs. Privacy attacks targeting the inference process can be broadly categorized into two types: (1) Output-based inversion attacks, which aim to reconstruct inputs from the model’s final outputs (e.g., probability distributions or embedding vectors), and (2) Intermediate state-based inversion attacks, which seek to extract information from the model’s internal computational states [15, 21, 34] (e.g., hidden states). Early work focused on inverting final representations [15], such as sentence embeddings, with methods like Vec2Text [21]. However, these approaches overlooked the richer information in deeper layers. The Embed Parrot attack [34] demonstrated that even deep hidden states can be exploited to accurately reconstruct user inputs, establishing the entire inference pipeline as a critical attack surface.
Unlike a hidden state (the fused output of a Transformer layer), the KV-cache contains the more primitive Key (K) and Value (V) vectors that serve as direct inputs to the attention calculation. Crucially, the KV-cache is a specialized data structure explicitly designed for persistence and reuse across generation steps to optimize inference. This unique role makes it a more potent and exploitable source of privacy leakage than other transient states.
Side-Channel Attacks Exploiting the KV-cache. Currently, research on attacks against the KV-cache has been primarily limited to side-channel attacks. A prominent example is PromptPeek [36], which exploits timing discrepancies created by KV-cache sharing in multi-tenant LLM services. An attacker observes variations in response latency to determine if their guessed prefix has hit another user’s cached entries, allowing them to reconstruct the prompt without direct data access.
However, the viability of such timing side-channel attacks is diminished by modern memory management optimizations. Contemporary inference engines, particularly those using PagedAttention [14], manage the KV-cache in non-contiguous memory blocks (pages), with a typical default block size of 16 tokens. This block-based management means that for a cache hit to occur, an attacker would need to correctly guess an entire block of 16 tokens simultaneously. For an LLM with a vocabulary size in the hundreds of thousands, the search space for a single block exceeds $100,000^{16}$, rendering the attack computationally infeasible. In contrast, our work considers a different, more direct threat model: if an attacker can gain direct access to the KV-cache content (e.g., through memory access or network eavesdropping), they can bypass these side-channel limitations. This direct access enables a more practical collision attack on a per-token basis, reducing the attack complexity to a feasible level and exposing a more fundamental vulnerability. Given that the KV-cache is commonly stored and transmitted in plaintext to preserve performance, this paper aims to systematically explore this underexplored problem of direct, content-based attacks.
III Attack Landscape: Inferring Private User Inputs from KV-cache
While KV-cache is crucial for enhancing the inference efficiency of LLMs, it concurrently introduces novel privacy vulnerabilities. The contextual information stored within KV-cache can contain private user data. Specifically, entries in the KV-cache often correspond directly to tokens in the user’s input sequence. An attacker gaining access to this KV-cache information could potentially infer the original user input, thereby compromising user privacy.
In this section, we aim to study an important research question: is an attacker able to conduct a prompt reconstruction attack to recover the user input from the KV-cache? We first present the threat model of the attack. Subsequently, we propose two distinct attacks to achieve this goal.
III-A Threat Model
In our threat model, we assume that an LLM server has deployed an open-source LLM based service. After a user uses the service with private user inputs, it generates a large number of KV-caches. The attacker gains access to the KV-caches and tries to infer the private inputs, as showed in Figure 2.
Figure 2: The attacker restores user input based on model parameters and leaked KV-cache.
Attacker’s Objective. The primary objective of the attacker is to recover the user input prompt from the accessed KV-cache. In this paper, we assume that the attacker attempts to restore the user input verbatim, which may significantly compromise the privacy of the user.
Attacker’s Capabilities. We consider an attacker with relatively strong yet still practical capabilities, as follows.
  • Access to KV-cache. The attacker can obtain the KV-caches stored by an LLM server. This can be achieved by eavesdropping on the communication channel between nodes (e.g. through the LeftoverLocals attack [28]), reading persisted KV-cache from storage, or other unspecified means of accessing the KV-cache pool.
  • Access to the foundation model on which the deployed LLM is developed. We assume a gray-box setting where the attacker can gain access to the foundation model’s weights on which the deployed LLM is developed [^1] This assumption is realistic given that now most LLM services are developed on open-source foundation models. First, disclosing the underlying foundation model is often a fundamental requirement of open-source licenses. Second, even if model information isn’t publicly disclosed, an attacker could identify the specific foundation model via model fingerprinting attacks [23, 7].. The deployed model could be a fine-tuned version of the foundation model.
III-B Input Reconstruction Attacks from KV-cache
This section investigates the risk of user input leakage from KV-cache data during LLM inference. We systematically present three attack vectors, namely inversion attack, collision attack, and injection attack, for reconstructing the original user inputs from the KV-cache. These attacks differ in complexity, applicability, and exploited vulnerabilities, collectively illustrating the threat landscape.
(a) KV-cache Inversion Attack.
(b) KV-cache Collision Attack.
(c) KV-cache Injection Attack.
Figure 3: Workflow of input reconstruction attacks.
III-B1 KV-cache Inversion Attack
The KV-cache Inversion Attack, as showed in Figure 3a, is a direct method that attempts to reverse-engineer the input to the attention layer by leveraging the KV-cache data in conjunction with the model’s known attention projection matrices. As defined in the forward computation (Eq. (1)), an attacker who obtains the KV-cache and knows the model’s attention layer parameters $W_{k}$ and $W_{v}$ could theoretically perform a reverse calculation to recover the attention layer’s input $x_{i}$, as Eq. (4).
$x_{i}=k_{i}({R}_{\Theta,i}^{d})^{-1}(W_{k}^{\top})^{-1},\quad x_{i}=v_{i}(W_{v}^{\top})^{-1}.$ (4)
However, the practicality of this direct inversion attack is severely constrained by two fundamental factors. First, it requires the attention projection matrices (e.g., $W_{k},W_{v}$) to be invertible. This condition is met by some earlier models using standard Multi-Head Attention (MHA), such as LLaMA-7B [32], but not by many contemporary models. Architectures like LLaMA 3 [9], Qwen [37], and DeepSeek [20] employ optimizations like Grouped-Query Attention (GQA)[3] or Multi-Head Latent Attention (MLA)[26], which often result in non-square, non-invertible projection matrices, rendering the attack impractical. Second, the attack’s effectiveness is largely confined to the first decoder layer’s KV-cache, as its inversion directly yields the input sequence’s embeddings. In contrast, inverting from deeper layers only recovers intermediate hidden states, from which reconstructing the original text is a significantly more complex and data-intensive problem often requiring auxiliary models [34].
Due to these significant limitations, the KV-cache Inversion Attack, while theoretically straightforward, has a narrow scope of applicability and does not pose a threat to most modern LLM architectures.
III-B2 KV-cache Collision Attack
Unlike the KV-cache Inversion Attack, the KV-cache Collision Attack operates by iteratively generating KV-cache entries for candidate tokens from the model’s vocabulary. For each candidate token appended to a partially reconstructed input sequence, the attack performs local model inference, generates the corresponding KV-cache, and compares it against the intercepted KV-cache to determine if the candidate token is correct for the current position.
Fundamental Principle of the Collision Attack. The collision attack reconstructs user input by identifying the token at each position that minimizes a distance metric (e.g., L2 distance) between a locally generated KV-cache and the leaked KV-cache. Its effectiveness hinges on the assumption that the distance for the correct token, $dis_{target}$, is statistically separable from the distances of all other incorrect tokens from the vocabulary, $dis_{other}$, allowing for its unambiguous identification.
General Attack Procedure. As illustrated in Figure 3b, the attack reconstructs the user’s input token-by-token. After initially intercepting the target KV-cache from the cloud service, the attacker enters an iterative loop for each unknown token position. First, guided by the already-reconstructed prefix, the attacker’s local model instance predicts a probabilistically-ranked list of candidate tokens. Next, for each high-probability candidate, the attacker performs a forward pass to generate a local KV-cache segment and compares it against the corresponding segment of the target KV-cache to test for a match (a “collision”). Upon a successful match, the correct token is confirmed and appended to the reconstructed prefix, and the attacker proceeds to the next position. This process continues until the entire input sequence is recovered.
Implementation Optimizations for the Collision Attack. A naive collision attack that processes the entire model vocabulary is impractical due to prohibitive VRAM consumption and execution time. We therefore introduce some synergistic optimizations to make the attack practical. First, instead of analyzing all prioritized candidates, we apply batched statistical validation, testing a subset of candidate tokens and computing the distribution of their distances ($dis_{batch}$) from the leaked KV-cache. The correct token is then identified as any distinct statistical outlier within this batch, under the assumption that the batch provides a representative sample of the distance distribution for incorrect tokens. Second, we employ a probability-guided search. Using the model’s own predictions based on the already reconstructed prefix, we prioritize collision tests for high-likelihood candidate tokens. Furthermore, we prune the search space by truncating the vocabulary and discarding tokens with very low probabilities. Together, these techniques reduce the memory footprint and bring the attack latency down to the order of seconds.
Enhancing Collision Attack with Prior Knowledge. A key challenge in the collision attack is setting an optimal classification threshold to distinguish the correct token’s distance ($dis_{target}$) from the distribution of incorrect token distances ($dis_{other}$). A basic attacker, lacking ground-truth data, must rely on a static, heuristic threshold, typically derived from the statistics of observed incorrect token distances (e.g., $\mu_{other}-3\sigma_{other}$). However, an adversary capable of performing chosen-plaintext attacks can generate KV-caches for known inputs. This prior knowledge enables a more precise statistical characterization of both the $dis_{target}$ and $dis_{other}$ distributions, allowing for a more effective, adaptive threshold that significantly boosts reconstruction accuracy.
The need for an adaptive threshold is formalized by the success probability for a token at rank $r$ in the probability-sorted vocabulary, as defined in Eq. (5):
$P(\text{success}|t)=P(dis_{other}>t)^{r-1}\times P(dis_{target}<t).$ (5)
This equation demonstrates that the optimal threshold $t$ is dependent on the token’s expected rank $r$, as it must balance the trade-off between correctly rejecting the r1r-1 preceding (incorrect) candidates and accepting the true token. A static heuristic, by contrast, cannot effectively manage this rank-dependent trade-off.
The KV-cache Collision Attack overcomes the limitations of the Inversion Attack concerning specific KV-cache and model attention mechanism requirements. An attacker who intercepts KV-cache from potentially any layer of various LLM architectures can, in principle, employ this collision-based approach to reconstruct user inputs. The practical effectiveness and performance of this attack will be demonstrated and evaluated in subsequent experimental sections.
III-B3 KV-cache Injection Attack
The KV-cache Injection Attack leverages the LLM’s own semantic capabilities to exfiltrate information, operating differently from direct state-matching methods. The KV-cache contains substantial information from user-model interactions, however, itself (comprising only $k$ and $v$ vectors for past tokens) lacks the $q$ vectors for the current generation step and thus cannot be directly used to continue model inference without a new input token. The core principle involves an attacker appending a crafted instruction (e.g., “Repeat the previous content.”) to an intercepted KV-cache, as showed in Figure 3c. This action provides the necessary query, key, and value vectors to resume inference from the compromised state, effectively tricking a standard LLM into executing the command within the context of the user’s private data. This can compel the model to “echo” or regenerate the sensitive input stored implicitly within the cache. While this method is generally less precise than the collision attack, its primary advantage is speed, requiring only a single forward inference pass. The viability of this attack vector reveals a crucial requirement for any comprehensive defense: a protected KV-cache must be rendered semantically unintelligible to prevent an unauthorized LLM from meaningfully parsing its contents or continuing generation from it.
We conduct comprehensive evaluations on these three attacks in Section V. The experimental results demonstrate the practicality and effectiveness of input reconstruction attacks from KV-cache, underscoring the urgent need for developing an effective defense technique.
IV KV-Cloak: A Defense Mechanism for
KV-cache Obfuscation
IV-A Motivation for KV-Cloak
Limitations of Existing Privacy-Preserving Techniques. While several techniques exist for data privacy, they prove impractical or inadequate for securing the LLM KV-cache, highlighting the need for a specialized solution. The introduction to existing privacy-preserving techniques and their limitations is as follows.
  • Cryptographic Methods: Standard cryptographic techniques [4, 31], such as symmetric AES encryption, provide confidentiality by ensuring data remains in ciphertext during storage and transmission, requiring decryption only for legitimate use. A more advanced approach, Homomorphic Encryption (HE) [22, 2], allows for direct computation on encrypted data, theoretically enabling parts of the attention mechanism to operate on the KV-cache without ever decrypting it. However, despite their strong security guarantees, the immense computational overhead and latency introduced by these methods are prohibitive. Given that the KV-cache can be tens or hundreds of gigabytes, applying full encryption or HE is unsuitable for the high-throughput, low-latency requirements of LLM inference.
  • Differential Privacy (DP): Differential Privacy is a rigorous model that protects individual privacy by injecting carefully calibrated random noise into data or query results [1, 16, 39]. The fundamental goal is to make the output of a computation statistically insensitive to the presence or absence of any single individual’s data in the input dataset. In the context of the KV-cache, this would involve adding noise to the Key and Value vectors. The primary limitation of this approach is the inherent trade-off between privacy and utility. To achieve a meaningful level of privacy, the required amount of noise would significantly degrade the LLM’s inference accuracy to an unacceptable degree.
  • KV-Shield: KV-Shield [38] is a lightweight obfuscation scheme and, to our best knowledge, the only privacy-preserving method designed specifically for the KV-cache. Its process involves synchronously permuting the rows of the model’s attention layer parameters ($W_{q},W_{k},W_{v}$). This operation effectively shuffles the element positions within the resulting KV-cache vectors, obfuscating them before an attention score is calculated. The obfuscated attention output is then de-obfuscated for subsequent steps. However, this approach has critical flaws. From a security perspective, its fixed shuffling pattern does not alter the statistical properties of the data, leaving it vulnerable to collision attacks. For compatibility, its technique is incompatible with modern LLM architectures that utilize features like Rotary Positional Embedding (RoPE), as the shuffling interferes with the position-dependent calculations.
Design Objectives. The shortcomings of existing methods reveal an urgent need for a defense mechanism that is both secure and practical. To be effective, such a solution must achieve three key objectives:
  • High Security: It must provide robust protection against targeted reconstruction attacks.
  • Minimal Accuracy Impact: It must operate without causing any degradation in the model’s output accuracy.
  • Low Performance Overhead: It must introduce only negligible latency into the inference pipeline.
To meet these challenges, we propose KV-Cloak, a novel defense mechanism designed to provide strong, lossless, and efficient protection for the KV-cache.
IV-B A Naive Defense: Obfuscation via Reversible Linear Transforms
We first evaluate a baseline defense that employs a reversible linear transform to obscure the statistical properties of the $k$ and $v$ vectors. In the context of modern inference frameworks like vLLM, we represent all key vectors for a single attention head within a KV-cache block as a matrix $K\in\mathbb{R}^{b\times d}$, where $b$ is the number of vectors per block and $d$ is the attention head dimension. As the obfuscation method is analogous for both keys and values, we use $K$ as our canonical example. The transformation is defined as:
$K^{\prime}=SKM,$ (6)
where $S\in\mathbb{R}^{b\times b}$ and $M\in\mathbb{R}^{d\times d}$ are secret, randomly generated, and invertible matrices. The objective is to transform the original matrix $K$ without altering its dimensions, preventing an adversary from directly recovering its contents.
Vulnerability Analysis. Despite its simplicity, this linear transform is fundamentally insecure under a Chosen-Plaintext Attack (CPA). While an adversary cannot craft an arbitrary matrix $K$ from scratch—as its contents are determined by the model’s embedding and projection layers—they can still exert significant influence through carefully constructed inputs. This capability is sufficient to break the scheme.
The fixed linear nature of the transformation makes it highly susceptible to algebraic cryptanalysis. An adversary can mount a differential attack by choosing two plaintexts, $K_{1}$ and $K_{2}$, thereby controlling the difference $\Delta K=K_{1}-K_{2}$ and observing the corresponding output difference $\Delta K^{\prime}=S(\Delta K)M$. By systematically choosing $\Delta K$ to be a series of standard basis matrices (i.e., matrices with a single non-zero entry), the adversary can isolate the columns of $S$ and the rows of $M$. This process allows for the full recovery of the secret matrices $S$ and $M$ (up to a trivial scaling ambiguity), leading to a complete compromise of the scheme.
IV-C The Improved Defense: One-Time Pad Block-wise Shuffling
Eliminating Redundant Positional Information in KV-cache. To enhance the security of the obfuscation scheme, we introduce additional randomness by eliminating redundant positional information inherent to KV-cache storage. During attention mechanism computations, the physical ordering of $k,v$ vectors in the KV-cache is an unnecessary redundancy for inference. In reality, positional information critical for computation is already embedded into the elements of $q$ and $k$ vectors through rotary positional encoding.
Figure 4: Shuffling $k,v$ vectors within a KV-cache block preserves attention computation results while disrupting positional order.
By randomly permuting $k,v$ within each block while maintaining their correspondence to input tokens $x$, we eliminate positional clues that attackers could exploit, and the results of the attention mechanism were not affected (as Figure 4). This permutation increases the computational complexity for attackers to match plaintext-ciphertext pairs by a factor of $b!$, where $b$ denotes the block size (default $b=16$ in vLLM).
The enhanced obfuscation process employs a one-time pad permutation matrix $\hat{P}$ to reorder vectors within blocks before applying invertible matrices $S$ and $M$:
$K^{\prime}=S\hat{P}KM.$ (7)
Crucially, $\hat{P}$ does not need to be stored as a key since the deobfuscated $\hat{P}K$ can directly participate in subsequent computations without reconstruction. This design choice preserves computational efficiency.
Additive Mask Matrices with Position-Identifiable Features. However, when the original $K$ matrix has rank $1$ (i.e., all row vectors are identical), the entropy of the permutation $\hat{P}$ will collapse. To prevent rank deficiency, KV-Cloak introduce an additive masking matrix $A$ before permutation. Thereby, we can ensure the transformed matrix $(K+A)$ maintains sufficient rank to preserve cryptographic strength of $\hat{P}$.
Since the one-time pad permutation matrix $\hat{P}$ is not stored, computing $\hat{P}A$ during inverse obfuscation becomes infeasible, and thus $\hat{P}K=\hat{P}(K+A)-\hat{P}A$ cannot be derived from $\hat{P}(K+A)$. To avoid storing the one-time pad permutation matrix $\hat{P}$, we exploit the sparsity and small-elements properties of $k$ and $v$ vectors to design the additive matrix $A$ with additional structural constraints. This enables us to infer $\hat{P}A$ based on the data characteristics of $\hat{P}(K+A)$.
Figure 5: Using an excerpt from “The Bitter Lesson” [30] (see Appendix -B) as the input of LLaMA-3.2-1b, the KV-cache data element distribution generated during the inference process.
As shown in Figure 5, we observe that the elements of $k$ and $v$ vectors typically remain below a small threshold $\theta_{K}$ (e.g., 100). This allows us to construct a matrix $A$ where positional information is embedded into $K+A$ via additive operations using values significantly larger than $\theta_{K}$ in specific row vectors. After applying row permutations, the positional identifiers in $\hat{P}(K+A)$ remain identifiable. During inverse obfuscation, the positional information of $\hat{P}A$ can be directly identified from $\hat{P}(K+A)$ by detecting these outliers. Subtracting $\hat{P}A$ from $\hat{P}(K+A)$ yields $\hat{P}K$ for subsequent inference, drastically reducing storage overhead for the one-time pad key.
The complete obfuscation transformation, as shown in Eq. (8), incorporates additional matrices $\hat{P}$ and $A$ alongside invertible linear transformations $S$ and $M$:
$K^{\prime}=S\hat{P}(K+A)M,$ (8)
[…]
IV-D Implicit Obfuscation via Operator Fusion
To mitigate the performance overhead introduced by the obfuscation and de-obfuscation processes during runtime, we now demonstrate how the obfuscation operators can be fused directly into the model’s weight matrices. This allows the primary cost of obfuscation to be paid offline, resulting in a minimal impact on online inference latency.
Fusing the Obfuscation Operator into Model Weights. We define three transformed vectors $(q^{m},k^{m},v^{m})$ and a transformed output projection matrix $W_{o}^{m}$ using two secret, invertible matrices, $M_{1}$ and $M_{2}$. Let $q,k,v$ be the original query, key, and value vectors, and let $W_{o}$ be the original output projection matrix. The transformations are defined as follows:
$\left\{\begin{array}[]{l}q^{m}=q(M_{1}^{-1})^{\top},\\ k^{m}=kM_{1},\\ v^{m}=vM_{2},\\ W_{o}^{m}=W_{o}(M_{2}^{-1})^{\top}.\end{array}\right$. (9)
These transformations are specifically designed to leave the core attention mechanism computations invariant. The attention scores remain unchanged, as shown by the dot product:
$q_{i}^{m}(k_{j}^{m})^{\top}$ $=\left(q_{i}(M_{1}^{-1})^{\top}\right)\left(k_{j}M_{1}\right)^{\top}=q_{i}(M_{1}^{-1})^{\top}M_{1}^{\top}k_{j}^{\top}$ (10)
$=q_{i}(M_{1}M_{1}^{-1})^{\top}k_{j}^{\top}=q_{i}k_{j}^{\top}.$
Similarly, the output of the attention head after projection remains identical:
$v_{j}^{m}(W_{o}^{m})^{\top}$ $=(v_{j}M_{2})\left(W_{o}(M_{2}^{-1})^{\top}\right)^{\top}$ (11)
$=v_{j}M_{2}(M_{2}^{-1})W_{o}^{\top}=v_{j}W_{o}^{\top}.$
This invariance guarantees that the model’s output is mathematically identical to that of the original, unprotected model, ensuring lossless accuracy.
Next, we incorporate the projection from the input embedding $x$ and the application of RoPE, denoted by the position-dependent matrix $R_{\Theta,i}^{d}$. The equations for the transformed vectors become:
$\left\{\begin{array}[]{l}q^{m}=xW_{q}^{\top}R_{\Theta,i}^{d}(M_{1}^{-1})^{\top}\\ k^{m}=xW_{k}^{\top}R_{\Theta,i}^{d}M_{1}\\ v^{m}=xW_{v}^{\top}M_{2}\\ W_{o}^{m}=W_{o}(M_{2}^{-1})^{\top}\end{array}\right.$. (12)
From these formulations, we can observe the path to operator fusion. The secret matrices $M_{2}$ and $M_{2}^{-1}$, which obfuscate the value vectors, can be straightforwardly fused into the model’s weights by pre-computing new weight matrices $(W_{v}^{m})^{\top}=W_{v}^{\top}M_{2}$ and $W_{o}^{m}=W_{o}(M_{2}^{-1})^{\top}$.
However, fusing $M_{1}$ and $M_{1}^{-1}$ into $W_{k}$ and $W_{q}$ is more complex due to the intermediate application of the position-dependent RoPE matrix $R_{\Theta,i}^{d}$. The fusion becomes feasible under a specific algebraic condition: if the secret matrix $M_{1}$ and the RoPE matrix $R_{\Theta,i}^{d}$ commute, i.e., $R_{\Theta,i}^{d}M_{1}=M_{1}R_{\Theta,i}^{d}$. If this condition holds, we can rewrite the expression for $k^{m}$:
$k^{m}=xW_{k}^{\top}R_{\Theta,i}^{d}M_{1}=x(W_{k}^{\top}M_{1})R_{\Theta,i}^{d}.$ (13)
This allows us to pre-compute a new, position-independent weight matrix $(W_{k}^{m})^{\top}=W_{k}^{\top}M_{1}$. A similar argument applies to fusing $(M_{1}^{-1})^{\top}$ into $W_{q}$.
Therefore, the viability of this operator fusion approach hinges on our ability to construct a secret matrix $M_{1}$ that commutes with the RoPE matrices. In Appendix -A, we formally prove that this commutativity is achieved when $M_{1}$ is structured as a block-diagonal matrix composed of $2\times 2$ rotation-scaling sub-matrices, a structure analogous to that of RoPE itself. Furthermore, to ensure the reversibility of the entire transformation for lossless de-obfuscation, $M_{1}$ must be constructed to be invertible.
By constraining $M_{2}$ to also be a random, invertible rotation-scaling matrix, and by carefully controlling the range of rotation-scaling coefficients in both $M_{1}$ and $M_{2}$, we ensure that the transformed padding tokens (which involve the affine transformation with matrix $A$) remain identifiable as outliers. These conditions allow the obfuscation operators ($M_{1},M_{2}$) and their corresponding de-obfuscation counterparts ($M_{1}^{-1},M_{2}^{-1}$) to be fused into the attention layer parameters $W_{q},W_{k},W_{v},\text{and }W_{o}$. Let the model parameters after operator fusion be denoted by $W_{q}^{m},W_{k}^{m},W_{v}^{m},W_{o}^{m}$. The new, fused weights are pre-computed as follows:
$\left\{\begin{aligned} W_{q}^{m}&=M_{1}^{-1}W_{q}\\ W_{k}^{m}&=M_{1}^{\top}W_{k}\\ W_{v}^{m}&=M_{2}^{\top}W_{v}\\ W_{o}^{m}&=W_{o}(M_{2}^{-1})^{\top}\end{aligned}\right.$. (14)
By deploying the LLM with these pre-computed fused weights, we effectively integrate the random invertible block-diagonal matrix $M_{1}$ and the general random invertible matrix $M_{2}\in\mathbb{R}^{d\times d}$ into the attention layer parameters. This eliminates the need for explicit online matrix multiplications with $M_{1}$ and $M_{2}$, significantly reducing performance overhead.
Ultimately, the KV-Cloak obfuscation formula, leveraging operator fusion, is applied as follows:
$K^{\prime}=S\hat{P}(K^{m}+A),$ (15)
where $K^{m}=KM$ represents the KV-cache generated by the inference process using the fused model weights. It is crucial to control the range of the rotation and scaling factors within the matrix $M$ to ensure that the additive mask $A$ can still reliably embed its positional information via the outlier mechanism.
TABLE I: Effectiveness of the inversion, collision, and injection attacks against the unprotected KV-cache from different model layers. Here, “collision+” denotes the collision attack enhanced with prior knowledge.
Model Metric Inversion Collision Collision+ Injection
First Mid Last First Mid Last First Mid Last All
LLaMA-7B BERTScore ($\uparrow$) 1.000 0.065 0.092 0.449 0.769 0.611 1.000 1.000 1.000 0.765
ROUGE-L ($\uparrow$) 1.000 0.036 0.062 0.500 0.562 0.436 1.000 1.000 1.000 0.687
LLaMA-3.2-1B BERTScore ($\uparrow$) 1.000 0.084 0.057 0.877 0.791 0.894 1.000 1.000 1.000 0.544
ROUGE-L ($\uparrow$) 0.994 0.038 0.000 0.709 0.617 0.680 0.994 0.994 0.994 0.315
LLaMA-3.2-3B-Instruct BERTScore ($\uparrow$) 0.055 0.095 0.083 0.782 0.668 0.820 1.000 1.000 1.000 0.540
ROUGE-L ($\uparrow$) 0.000 0.000 0.000 0.732 0.456 0.621 0.994 0.994 0.994 0.324
LLaMA-3.1-8B BERTScore ($\uparrow$) 0.071 0.061 0.062 0.873 0.652 0.764 1.000 1.000 1.000 0.616
ROUGE-L ($\uparrow$) 0.000 0.000 0.001 0.825 0.443 0.564 0.994 0.994 0.994 0.447
LLaMA-3.1-8B-Distilled BERTScore ($\uparrow$) 0.083 0.062 0.081 0.642 0.492 0.635 0.894 0.258 0.762 0.610
ROUGE-L ($\uparrow$) 0.000 0.000 0.001 0.633 0.227 0.413 0.868 0.122 0.479 0.421
Qwen2.5-Math-7B BERTScore ($\uparrow$) 0.229 0.105 0.105 0.918 0.552 0.783 1.000 0.983 0.996 0.422
ROUGE-L ($\uparrow$) 0.186 0.000 0.000 0.842 0.355 0.580 1.000 0.977 0.996 0.286
With this optimization, the online computational overhead is drastically reduced. The matrix multiplication with $M$ is now part of the offline weight preparation. The online cost for obfuscating and de-obfuscating one KV-cache block is now dominated by the multiplications with $S,\hat{P},S^{-1}$, totaling approximately $b^{3}+2b^{2}d$ floating-point operations. For the LLaMA-3.1-8B example ($b=16,d=128,D=4096$), this overhead is merely $0.83\%$ of the cost of re-computing the KV-cache block, rendering the performance impact minimal. Detailed analysis can be found in Appendix -F.
V Evaluation
V-A Experimental Setup
Models. We first select several state-of-the-art models from the popular LLaMA [9] and Qwen [37] series, including LLaMA-3.2-1B, LLaMA-3.2-3B-Instruct, LLaMA-3.1-8B, Qwen2.5-Math-7B, and DeepSeek-R1-Distill-LLaMA-8B (LLaMA-3.1-8B-Distilled) [10], which is fine-tuned from LLaMA-3.1-8B. In addition to these five models that use Grouped-Query Attention (GQA), we also leverage a classic models with Multi-Head Attention (MHA), namely LLaMA-7B [32], to demonstrate the generalizability of our method.
Datasets. To evaluate the effectiveness of our attacks, we use the lmsys-chat-1m dataset [43], which is collected from a real-world inference service and contains a rich variety of dialogues and instructions. We randomly sample 1,000 instances from this dataset as the test set for validating attack effectiveness. To assess the impact of KV-Cloak on model accuracy and inference latency, we employ standard language model benchmarks, including MMLU [13, 12] and SQuAD [25].
Attack and Defense Evaluation Metrics. BERTScore [41] and ROUGE-L [19] are used to measure the similarity between the original input and the text reconstructed from the KV-cache. BERTScore, which is based on the all-mpnet-base-v2 model, is better at capturing semantic similarity, while ROUGE-L primarily reflects lexical-level precision and recall.
V-B Attack Effectiveness
In this section, we first evaluate the privacy leakage risk of using the KV-cache by implementing our proposed inversion, collision, and injection attacks. Additional experiments, such as the ablation study of attacks, can be found in Appendix -C.
Experimental Settings. We attack 7 LLMs as described in Section V-A. Notably, while attacking the LLaMA-3.1-8B-Distilled, which is a fine-tuned version, we assume that the attacker can only get access to its base model (i.e., LLaMA-3.1-8B). It can be utilized to validate the attack effectiveness against fine-tuning models. Furthermore, the inversion and collision attacks require KV-cache data from only a single layer. Given that models have varying numbers of layers, we select the KV-cache from the first, middle, and last layers to perform the inversion and collision attacks. In contrast, the injection attack is conducted on the complete KV-cache data, encompassing all layers.
Results of KV-cache Inversion Attack. As shown in Table I, our experimental results generally align with our analysis: a successful inversion attack requires two sufficient conditions to be met: (1) access to the KV-cache from the model’s first layer, and (2) an invertible attention projection matrix, which is characteristic of the MHA mechanism. The inversion attack can reconstruct the user input with near-perfect fidelity (approaching 100%) from the first-layer KV-cache of MHA-based models. In contrast, reconstruction attempts using the KV-cache from subsequent layers yield completely unintelligible results (mostly <0.1<0.1).
However, we also find that for the LLaMA-3.2-1B model using GQA, the inversion attack achieved a reconstruction fidelity of nearly 100% on the first-layer KV-cache. This may be because the ratio of the rank of this model’s attention projection matrix to its hidden state dimension is particularly high. Consequently, even though the projection matrix is not a square, invertible matrix, the least squares method can still recover the input to the attention layer with high accuracy.
Takeaway 1: Inversion attack is mostly effective against the KV-cache of the first MHA layer.
Results of KV-cache Collision Attack. For the collision attack, we measure the distance between a locally generated KV-cache and the leaked KV-cache using the Frobenius norm. We experimentally observed that the distances ($dis_{other}$) between the leaked KV-cache and the caches generated from incorrect tokens in the vocabulary approximate a Gaussian distribution, as shown in Figure 8. Consequently, we leverage the statistical properties of this distribution (i.e., its mean and standard deviation) to identify the correct token, whose corresponding distance ($dis_{target}$) will manifest as a statistical outlier.
After experimental analysis, we defined outliers as values falling below 3 standard deviations from the mean of the $dis_{batch}$ distribution. The batch size for local analysis was set to 256. Details of the ablation study can be found in Appendix -C1. Under these settings, the experimental results for the collision attack are presented in the “Collision” column of Table I, demonstrating high reconstruction accuracy across all layers of all models.
We also experimented with truncating the probability-sorted vocabulary on the LLaMA-3.2-1B model, aiming to reduce the time lost on unlikely tokens. The results are presented in Figure 6. We found that even when searching only the top $1/8$ of tokens with the highest predicted probabilities, the average input reconstruction fidelity remains at approximately 96.1% of what is achieved by searching the entire vocabulary. Meanwhile, the attack time is reduced to less than 42.9% of the full search, with the average time to reconstruct a user input from a single layer’s KV-cache dropping from 5.06s to 2.17s. This confirms that truncating the vocabulary based on model-predicted probabilities is a highly effective strategy for accelerating the collision attack.
Figure 6: The effect of truncating the probability-sorted vocabulary on reconstruction fidelity and attack time. Experiments were run with a batch size of 256 and an outlier threshold of $3\sigma_{other}$.
The experimental results for the collision attack demonstrate that this method overcomes the limitations of the inversion attack, which is restricted to specific layers (i.e., the first layer) and attention mechanisms (i.e., MHA). For open-source models, the collision attack can reconstruct user inputs with high accuracy and acceptable efficiency from the KV-cache of any layer. Crucially, it is also effective against the LLaMA-3.1-8B-Distilled model—a model fine-tuned from Meta-LLaMA-3.1-8B—by using the public weights of its open-source base model.
Takeaway 2: Collision attack achieves relatively high accuracy, demonstrating its universality and the high-risk threat it poses to real-world systems.
Collision Attack Enhanced with Prior Knowledge (Collision+). As shown in Table I, our enhanced collision attack (assuming rank $r=8$) is highly potent, achieving near-perfect reconstruction accuracy (approaching 100%) on any layer of any tested open-source model, while also demonstrating notable effectiveness against fine-tuned models.
We also conducted experiments on the LLaMA-3.1-8B-Distilled model using optimal thresholds derived for different assumed ranks ($r$). The results are shown in Figure 7. For the first layer, the actual average rank of the target token is approximately 4; using the optimal threshold for this rank increases the reconstruction accuracy to 138.6% of the unenhanced attack. For the middle layer, the actual rank appears to be greater than 128, as none of the tested enhanced thresholds surpassed the accuracy of the baseline attack. For the last layer, the optimal rank is between 64 and 128, and applying the corresponding enhanced threshold boosts the reconstruction accuracy to approximately 139.0% of the baseline.
Figure 7: Collision attack experiments on LLaMA-3.1-8B-Distilled using optimal thresholds derived for different assumed token ranks ($r$).
Takeaway 3: Augmenting the collision attack with model-specific prior knowledge achieves near-perfect ($\approx 100\%$) user-input recovery accuracy.
(a) The distributions of LLaMA-3.1-8B model.
(b) The distributions of LLaMA-3.1-8B-Distilled model.
Figure 8: Distributions of $dis_{target}$ (orange) and $dis_{other}$ (blue) from a collision attack. The input is an excerpt from “The Bitter Lesson”. The attack uses the LLaMA-3.1-8B model’s weights against the last-layer KV-cache generated by itself (left) and by LLaMA-3.1-8B-Distilled (right). The green dotted line shows the baseline heuristic threshold ($3\sigma_{other}$), while the red dashed line shows the enhanced threshold derived from prior knowledge (assuming rank $r=64$).
To further illustrate this, we analyzed the distance distributions using an excerpt from “The Bitter Lesson” as input. Figure 8 shows these distributions for an attack on the last-layer KV-cache. We observe that for the open-source model (left panel), the baseline heuristic threshold of $3\sigma_{other}$ results in a 0.13% false-positive rate (i.e., 0.13% of $dis_{other}$ distances are misclassified). If we assume the correct token is at rank $r=64$ in the sorted vocabulary, this yields a per-token attack success rate of 91.84%. In contrast, by leveraging a more precise threshold derived from statistical analysis, this success rate can be boosted to a perfect 100%. For the fine-tuned model (right panel), the enhanced threshold improves the per-token success rate from 91.79% to 97.82%. This substantial improvement in per-token accuracy leads to a dramatic increase in the overall reconstruction fidelity for the entire input sequence.
Results of KV-cache Injection Attack. The success of this attack hinges on two factors: the model’s ability to effectively parse the semantic content of the KV-cache and its comprehension of the injected instruction. For our evaluation, we employed the most effective instruction identified in our ablation study, “Repeat the previous content.” (the selection process is detailed in Appendix -C2), to attack the KV-cache data from each target model. As shown in Figure I, the injection attack yielded an average BERTScore of 0.58 and a ROUGE-L score of 0.42 for the reconstructed user inputs. These results are consistently lower than those achieved by the more direct collision attack.
The attack performed best on the LLaMA-7B. This supports the hypothesis that the MHA mechanism preserves more contextual information than the more compressed attention variants (e.g., GQA) within the KV-cache.
Takeaway 4: Absent verbatim recovery, injection attacks can still exfiltrate the core information of user inputs through semantic-level interaction.
The reconstruction accuracy of the injection attack is generally lower than that of the collision attack, and it requires access to the complete KV-cache from all layers. However, its key advantage is speed: the attack’s execution time is equivalent to a single inference step, making it significantly faster than the iterative collision attack. Moreover, the existence of this attack vector imposes an additional requirement on any defense mechanism: it must ensure that a protected KV-cache is rendered unintelligible even to a standard, unprotected LLM.
V-C Evaluation of KV-Cloak
V-C1 Experimental Settings
In order to comprehensively evaluate our proposed KV-Cloak scheme, this section presents a systematic comparison against two baselines: (1) a standard, unprotected system (Plaintext) and (2) a defense based on differential privacy with Gaussian noise (DP). Our evaluation is structured around three core dimensions: Security, Model Accuracy, and Performance Overhead. The results demonstrate that KV-Cloak is the only solution that provides robust security protection with acceptable performance overhead, all without sacrificing model accuracy.
In our KV-Cloak, the obfuscation involves matrix multiplication of KV-cache blocks with invertible secret matrices $S$ and $M$, along with the addition of a random mask $A$.
  • Block Size: To ensure compatibility with the PagedAttention mechanism used in modern inference engines like vLLM, we experimented with common block sizes: 16, 32, and 64.
  • Invertible Matrices $S,M$: To avoid precision loss from matrix inversion during de-obfuscation, we sample $S$ and $M$ from the space of orthogonal matrices.
  • Additive Mask $A$ and Padding: The values in the mask $A$ and the padding for incomplete blocks are kept small to minimize their impact on numerical precision. We sample the elements of $A$ from the range $3\theta_{K},4\theta_{K}]$ and use a value of $1.5\theta_{K}$ for padding. Here, $\theta_{K}$ is the maximum absolute value observed in the K-cache elements when inferring on an excerpt from “The Bitter Lesson”. The same methodology is applied for the V-cache.
To establish a robust DP baseline, we conducted an ablation study (detailed in Appendix -D) to select a configuration that balances utility and privacy for comparison against KV-Cloak. Based on our findings, we selected a practical configuration: the clipping threshold was set to the 50th percentile of the L2 norm distribution observed across the dataset. Subsequently, we applied Gaussian noise calibrated to satisfy ($\epsilon=10^{8},\delta=10^{-5}$)-DP independently to both the K-cache and V-cache.
V-C2 Evaluation of Security
TABLE II: Efficacy of Defense Mechanisms Against Input Reconstruction Attacks.
Model Protect Type Metric Inversion Collision Collision+ Injection
First First Mid Last First Mid Last All
LLaMA-7B Original BERTScore ($\downarrow$) 1.000 0.449 0.769 0.611 1.000 1.000 1.000 0.765
ROUGE-L ($\downarrow$) 1.000 0.500 0.562 0.436 1.000 1.000 1.000 0.687
KV-Cloak BERTScore ($\downarrow$) 0.091 0.070 0.069 0.071 0.036 0.036 0.036 0.082
ROUGE-L ($\downarrow$) 0.068 0.000 0.000 0.000 0.044 0.044 0.044 0.000
($10^{8},10^{-5}$)-DP BERTScore ($\downarrow$) 0.085 0.082 0.672 0.344 0.109 0.937 0.991 0.085
ROUGE-L ($\downarrow$) 0.065 0.041 0.433 0.197 0.097 0.901 0.979 0.009
LLaMA-3.2-1B Original BERTScore ($\downarrow$) 1.000 0.877 0.791 0.894 1.000 1.000 1.000 0.544
ROUGE-L ($\downarrow$) 0.994 0.709 0.617 0.680 0.994 0.994 0.994 0.315
KV-Cloak BERTScore ($\downarrow$) 0.085 0.072 0.074 0.069 0.051 0.051 0.051 0.079
ROUGE-L ($\downarrow$) 0.009 0.000 0.000 0.000 0.002 0.002 0.002 0.000
($10^{8},10^{-5}$)-DP BERTScore ($\downarrow$) 0.633 0.849 0.763 0.849 0.973 0.995 1.000 0.393
ROUGE-L ($\downarrow$) 0.622 0.604 0.587 0.604 0.966 0.989 0.994 0.248
LLaMA-3.1-8B-Distilled Original BERTScore ($\downarrow$) 0.083 0.642 0.492 0.635 0.885 0.251 0.829 0.610
ROUGE-L ($\downarrow$) 0.000 0.633 0.227 0.413 0.858 0.112 0.552 0.421
KV-Cloak BERTScore ($\downarrow$) 0.093 0.070 0.070 0.069 0.041 0.041 0.041 0.088
ROUGE-L ($\downarrow$) 0.002 0.000 0.000 0.000 0.003 0.003 0.003 0.000
($10^{8},10^{-5}$)-DP BERTScore ($\downarrow$) 0.079 0.320 0.440 0.566 0.526 0.267 0.824 0.118
ROUGE-L ($\downarrow$) 0.003 0.291 0.185 0.351 0.530 0.122 0.543 0.049
This section evaluates the capability of different protection schemes to defend against our three KV-cache-based reconstruction attacks. We repeat the attack experiments on the KV-cache after applying each defense mechanism. As shown in Table II (evaluation on the remaining models can be found in Appendix -E), for the “Plaintext” baseline, all three attacks achieve very high success rates. After applying KV-Cloak, the semantic similarity (BERTScore) of all attack outputs drops to a level consistent with random chance, and the ROUGE-L score falls to nearly 0. These results are statistically indistinguishable from comparing the original input with a random string, demonstrating that semantic reconstruction is entirely disrupted. This proves that KV-Cloak effectively protects the private information within the KV-cache. For the DP baseline, its defensive efficacy is strongly correlated with the privacy budget $\epsilon$. With a weak budget of $\epsilon=10^{8}$, the accuracy of the inversion and injection attacks is reduced, but the collision attack can still recover some useful information.
We also compare the distributions of $dist_{target}$ and $dist_{other}$ for the collision attack on the last layer of LLaMA-3.2-1B’s KV-cache under different protections. As shown in Figure 9, the two distributions remain effectively distinguishable under DP protection. When using an enhanced threshold derived from prior knowledge, the per-token attack success rate is as high as 94.84% for ($10^{7},10^{-5}$)-DP, and achieving nearly 100% (the same as plaintext) for ($10^{7},10^{-5}$)-DP. In contrast, under KV-Cloak’s protection, the two distributions become completely indistinguishable, resulting in a 0% attack success rate.
(a) The distributions of LLaMA-3.2-1B model without protection.
(b) The distributions of LLaMA-3.2-1B model with KV-Cloak.
(c) The distributions of LLaMA-3.2-1B model with ($10^{7},10^{-5}$)-DP.
(d) The distributions of LLaMA-3.2-1B model with ($10^{8},10^{-5}$)-DP.
Figure 9: Distributions of $dis_{target}$ (orange) and $dis_{other}$ (blue) for a collision attack on the last-layer KV-cache. The input is an excerpt from “The Bitter Lesson” to the LLaMA-3.2-1B model. The distributions are shown under four conditions: (a) Plaintext, (b) KV-Cloak protection, (c) ($10^{7},10^{-5}$)-DP protection, and (d) ($10^{8},10^{-5}$)-DP protection. The green dotted line is the baseline heuristic threshold ($3\sigma_{other}$), and the red dashed line is the enhanced threshold assuming rank $r=8$.
To assess the robustness of our defense, we also tested KV-Cloak’s performance against the enhanced collision attack on multiple LLMs. As illustrated in Table II, our mechanism demonstrates strong defensive capabilities. With KV-Cloak enabled, the reconstruction accuracy of the attack drops to near-zero, and the recovered outputs are qualitatively equivalent to random noise, confirming that KV-Cloak effectively neutralizes this advanced attack vector.
Takeaway 5: KV-Cloak completely thwarts all our proposed attacks, reducing the quality of any reconstructed text to a level statistically indistinguishable from random noise.
V-C3 Inference Accuracy
To evaluate the model fidelity of our KV-Cloak, we simulate an inference service with a prefill-decode architecture. The KV-cache generated during the prefill phase is protected using either DP or KV-Cloak. This protected cache is then passed to the decode node (after de-obfuscation in KV-Cloak’s case) for subsequent token generation. We then evaluate this protected inference service on the MMLU and SQuAD benchmarks to measure the impact of each method on model accuracy.
TABLE III: Impact of KV-Cloak on inference accuracy (higher is better) across various models, using a block size of 16.
Model Plaintext KV-Cloak ($10^{8},10^{-5}$)-DP
MMLU SQuAD MMLU SQuAD MMLU SQuAD
LLaMA-7B 0.304 0.646 0.304 0.652 0.016 0.000
LLaMA-3.2-1B 0.335 0.457 0.335 0.458 0.262 0.258
LLaMA-3.2-3B-Instruct 0.619 0.652 0.619 0.652 0.379 0.012
LLaMA-3.1-8B 0.668 0.708 0.668 0.709 0.283 0.026
LLaMA-3.1-8B-Distilled 0.584 0.568 0.584 0.570 0.108 0.001
Qwen2.5-Math-7B 0.620 0.630 0.620 0.632 0.042 0.000
We tested the impact of KV-Cloak on all experimental models. The results in Table III show that across both benchmarks, the impact of KV-Cloak on inference accuracy is negligible. This confirms our theoretical design: KV-Cloak is a lossless protection scheme that does not degrade the model’s generation quality or its performance on downstream tasks.
Takeaway 6: KV-Cloak is virtually lossless, preserving the model’s fidelity and core utility without any degradation.
V-C4 Evaluation of Computational Overhead
Integrating a new cryptographic primitive into a highly-optimized system like vLLM presents a significant engineering challenge. Therefore, to accurately assess the performance overhead of KV-Cloak and demonstrate its compatibility with PagedAttention, we employ a dual approach: direct latency measurement combined with architectural analysis.
To quantify the computational overhead of KV-Cloak, we measured the inference time for the MMLU benchmark under its protection. We selected this benchmark specifically to evaluate a worst-case scenario: the MMLU task requires applying KV-Cloak’s obfuscation and de-obfuscation to the entire KV-cache of a long input sequence, yet generates only a single output token. This setup maximizes the ratio of cryptographic overhead to inference work. Furthermore, our measurements are conservative as they are based on a serial PyTorch-level implementation without any custom CUDA kernel optimizations, and they do not factor in network communication latency.
TABLE IV: Performance overhead of KV-Cloak (with a block size of 16) on various models during the MMLU benchmark.
Model Inference Time(s)
Plaintext KV-Cloak
LLaMA-7B 10193.1 10649.0 +4.47%
LLaMA-3.2-1B 1956.8 2155.6 +10.16%
LLaMA-3.2-3B-Instruct 4525.7 4821.4 +6.53%
LLaMA-3.1-8B 9797.4 10135.0 +3.45%
LLaMA-3.1-8B-Distilled 9832.8 10297.7 +4.73%
Qwen2.5-Math-7B 9184.7 9452.3 +2.91%
The results of our evaluation across various models are presented in Table IV. The findings show that KV-Cloak introduces a modest average latency overhead of approximately 5%. Notably, we observe a favorable scaling property: the relative overhead decreases as the model’s hidden size increases. In a real-world distributed deployment, the inclusion of network latency between prefill and decode nodes would make the relative impact of KV-Cloak’s computational overhead on the total end-to-end time even lower. These empirical results align with our theoretical analysis, confirming that with standard production optimizations, such as operator fusion, the performance impact of KV-Cloak is expected to be minimal.
Takeaway 7: KV-Cloak introduces negligible computational overhead to the inference pipeline.
VI Conclusion
This research exposes a critical security flaw at the heart of modern LLM inference systems: the privacy risk of data leakage from the KV-cache. We have demonstrated the feasibility of reconstructing sensitive user inputs through three novel attack strategies, with our collision attack proving particularly effective across various models. This underscores the urgent need for dedicated protection mechanisms that do not compromise the efficiency gains the KV-cache provides.
In response, we designed KV-Cloak. By employing a lightweight, reversible obfuscation technique, KV-Cloak neutralizes the identified threats without degrading model accuracy or imposing significant latency. It is designed for seamless integration into existing high-performance inference frameworks like vLLM. Our work provides a vital contribution to building secure and trustworthy AI, offering a blueprint for balancing the competing demands of performance and user privacy in the next generation of LLM services. It establishes that strong privacy protection can be achieved without sacrificing the utility and efficiency that have made these models so powerful.
Ethic Considerations
This research aims to enhance the privacy and trustworthiness of LLM inference, but we acknowledge the dual-use nature of the vulnerabilities and attack methods we have uncovered. To fulfill our ethical responsibilities and mitigate any potential for misuse, we have committed to a policy of responsible disclosure. Prior to the public dissemination of this paper, we will share our findings, including the details of the vulnerabilities and our proposed defense, with the developers of major affected inference frameworks, such as vLLM. Furthermore, all of our attack validation experiments were conducted exclusively on public and non-sensitive academic datasets. No real user data was involved at any stage of our research. We firmly believe that by taking these responsible measures, the positive contributions of our defensive work, KV-Cloak, will far outweigh the risks associated with the disclosure of these attacks. We are confident that this work will encourage the community to build more trustworthy AI services that are secure by default.

### v3

@@
 However, this efficiency optimization introduces significant yet underexplored privacy risks.
-This paper provides the first comprehensive analysis of these vulnerabilities, demonstrating that an attacker can reconstruct sensitive user inputs directly from the KV-cache.
+This paper provides the first comprehensive analysis of these vulnerabilities, demonstrating that an adversary can reconstruct sensitive user inputs directly from the KV-cache.
 We design and implement three distinct attack vectors: a direct Inversion Attack, a more broadly applicable and potent Collision Attack, and a semantic-based Injection Attack.
@@
 This dramatically accelerates inference speed and improves throughput, especially for tasks involving long contexts or interactive sessions, making LLMs more practical for real-world deployment.
-Figure 1: LLM inference service workflow in a privacy-preserving setting and its associated security risks.
+Figure 1: Overview of the privacy-preserving LLM inference workflow and the associated KV-cache leakage threat model.
+While user-server communication is encrypted (black lines), the KV-cache is often transmitted and stored in plaintext (red dashed lines), creating a surface for privacy attacks.
 However, the storage and potential sharing of the KV-cache introduce significant yet underexplored privacy concerns [41, 39], as illustrated in Figure 1.
-This vulnerability stems from a critical trade-off made in production systems: while end-user communication with the LLM service is typically encrypted (represented by the black lines in our figure), the KV-cache itself is almost always processed, transmitted between compute nodes, and persisted in plaintext.
+This vulnerability stems from a critical trade-off made in production systems: while end-user communication with the LLM service is typically encrypted (represented by the black lines), the KV-cache itself is almost always processed, transmitted between compute nodes, and persisted in plaintext.
 This design choice is a concession to performance, as the unacceptable latency overhead from cryptographically securing the often gigabyte-scale KV-cache would violate the stringent demands of real-time inference.
-The feasibility of exploiting such plaintext data exposure is not merely theoretical.
-It was recently demonstrated by the LeftoverLocals attack [28].
-This vulnerability in certain AMD GPUs enables an unprivileged attacker to read data remnants from the GPU’s shared on-chip memory.
-This provides a concrete method to intercept and exfiltrate plaintext KV-cache data originating from other users’ concurrent inference sessions, proving that sensitive conversational data is indeed at risk.
-Crucially, the KV-cache is inherently derived from and directly correlated with user inputs.
-Consequently, if an attacker gains access to the KV-cache, they could potentially reverse-engineer sensitive information in the user’s original prompts or private data [6, 17].
-A critical research gap exists in understanding and mitigating the privacy risks associated with KV-caches in LLMs, necessitating further in-depth research.
+Notably, this risk is amplified in the emerging paradigm of confidential Model-as-a-Service (MaaS).
+Here, TEE adoption for inference privacy [5, 10] introduces a critical architectural trade-off: high-throughput architectures require deliberately externalizing the massive KV-cache from the TEE’s protection boundary.
+This architectural design—not a vulnerability—directly exposes the KV-cache to the Cloud Service Provider (CSP), who inherently possesses it without requiring any additional exploit [43].
+Crucially, the direct correlation between the KV-cache and user inputs [7, 21], combined with the architectural realities that expose it in plaintext, creates a severe and practical attack vector.
+A critical research gap therefore exists in mitigating these risks, necessitating a practical and secure defense mechanism.
 In this paper, to bridge the research gap, we present the first comprehensive study on the privacy risks of KV-cache in LLM inference.
 Specifically, we primarily study and answer the following two research questions:
-  • RQ1: Is an attacker able to reconstruct user inputs from the KV-caches?
+  • RQ1: Is an adversary able to reconstruct user inputs from the KV-caches?
 To address this question, we conduct a systematic investigation, demonstrating that attacks against the KV-cache are not only feasible but also diverse and broadly applicable.
@@
 The second, more general-purpose approach is the KV-cache Collision Attack.
-This attack reframes input reconstruction as a matching task based on forward computation: an attacker iteratively uses a local model instance to generate KV-caches for candidate inputs and compares them against the intercepted target KV-cache to find a match.
+This attack reframes input reconstruction as a matching task based on forward computation: an adversary iteratively uses a local model instance to generate KV-caches for candidate inputs and compares them against the intercepted target KV-cache to find a match.
 Because this method does not rely on any reverse computation, it has broader applicability.
@@
 The core of this attack is to leverage the LLM’s powerful capability to understand and execute instructions.
-By appending a specific instruction, such as “Repeat the previous content,” to the end of an intercepted KV-cache context, an attacker can induce the model to “echo” or generalize the core information held within the KV-cache.
+By appending a specific instruction, such as “Repeat the previous content,” to the end of an intercepted KV-cache context, an adversary can induce the model to “echo” or generalize the core information held within the KV-cache.
 Collectively, these attacks reveal that privacy leakage from the KV-cache is not merely a theoretical concern.
@@
 Even specialized solutions like KV-Shield [41], while lightweight, suffer from critical security flaws.
-Their fixed shuffling mechanism, as we analyzed in Section IV-A, is vulnerable to statistical analysis and is incompatible with modern LLM architectures that use features like Rotary Positional Embedding (RoPE).
-These limitations highlight the urgent need for a novel defense mechanism, .
+Their fixed shuffling mechanism, as we analyzed in Section V-A, is vulnerable to statistical analysis and is incompatible with modern LLM architectures that use features like Rotary Positional Embedding (RoPE)[32].
+These limitations highlight the urgent need for a novel defense mechanism.
 Therefore, we propose KV-Cloak, a lightweight and secure mechanism for KV-cache obfuscation designed to overcome these challenges.
@@
 Its security is multi-layered: it applies secret invertible linear transformations to obscure statistical properties, and critically, introduces a one-time random permutation matrix for each data block.
-This dynamic permutation prevents attackers from building stable algebraic relations across multiple queries.
+This dynamic permutation prevents adversaries from building stable algebraic relations across multiple queries.
 To further enhance performance, KV-Cloak utilizes operator fusion, an optimization where a portion of the secret obfuscation matrices is algebraically fused into the LLM’s attention layer weights offline.
@@
 We further prove that our proposed KV-Cloak is a highly practical solution that achieves robust security, near-lossless model fidelity, and high efficiency simultaneously.
-Our experiments show that KV-Cloak, with its negligible impact on accuracy and minimal latency overhead (mostly $10\%$), successfully resolves the stark trade-off between security and utility that plagues alternative approaches such as DP.
+Our experiments show that KV-Cloak, with its negligible impact on accuracy and minimal latency overhead (mostly $1\%$), successfully resolves the stark trade-off between security and utility that plagues alternative approaches such as DP.
 II Background and Related Work
 II-A Transformer-based LLM Inference
-Prevailing large language models, such as LLaMA [32], DeepSeek [20] and Qwen [37], are predominantly based on the Transformer decoder architecture [42].
+Prevailing LLMs, such as LLaMA [35], DeepSeek [24] and Qwen [40], are predominantly based on the Transformer decoder architecture [47].
 Self-Attention Mechanism.
-The core of the Transformer architecture is the self-attention mechanism.
-It captures dependencies between different positions within an input sequence, allowing the model to flexibly focus on various parts of the sequence to better understand the context.
-For an input sequence $X=(x_{1},\ldots,x_{n})$, the self-attention layer first applies linear transformations to $x_{i}$ at each position to obtain Query ($q$), Key ($k$), and Value ($v$) vectors,
-$q_{i}=x_{i}W_{q}^{\top}{R}_{\Theta,i}^{d},\quad k_{i}=x_{i}W_{k}^{\top}{R}_{\Theta,i}^{d},\quad v_{i}=x_{i}W_{v}^{\top}.$ (1)
-Here, $W_{q},W_{k},W_{v}$ are learnable weight matrices for the attention layer, and ${R}_{\Theta,i}^{d}$ denotes the Rotary Position Embedding (RoPE)[29].
-Then, the self-attention layer computes the dot product of the current query $q_{i}$ with all preceding keys $k_{j}$.
-The results are scaled, and a softmax function is applied to obtain attention weights $a_{ij}$.
-These weights are then used to compute a weighted sum of the value vectors.
-Finally, this sum is passed through a linear projection layer $W_{o}$ to produce the output $o_{i}$ for the current position.
-This process is detailed in Eq.
-(2), where $d$ is the dimension of each attention head.
-$a_{ij}=\frac{\exp\left(q_{i}{k_{j}}^{\top}/\sqrt{d}\right)}{\sum_{t=1}^{i}\exp\left(q_{i}{k_{t}}^{\top}/\sqrt{d}\right)},o_{i}=\left(\sum_{j=1}^{i}a_{ij}v_{j}\right)W_{o}^{\top}.$ (2)
-This mechanism enables the model to dynamically weigh the importance of different parts of the input sequence, thereby capturing complex contextual dependencies.
+For an input token $x_{i}$, the self-attention layer generates Query ($q_{i}$), Key ($k_{i}$), and Value ($v_{i}$) vectors via linear transformations.
+Crucially, modern architectures incorporate RoPE, denoted as ${R}_{\Theta,i}^{d}$:
+$q_{i}=x_{i}W_{q}^{\top}{R}_{\Theta,i}^{d},\;\;k_{i}=x_{i}W_{k}^{\top}{R}_{\Theta,i}^{d},\;\;v_{i}=x_{i}W_{v}^{\top}.$ (1)
+where $W_{(\cdot)}$ are learnable weight matrices.
+The attention score $a_{ij}$ is computed as the scaled dot product between $q_{i}$ and preceding keys $k_{j\leq i}$.
+The final output $o_{i}$ is the weighted sum of values projected by $W_{o}$:
+$a_{ij}=\frac{\exp(q_{i}{k_{j}}^{\top}/\sqrt{d})}{\sum_{t=1}^{i}\exp(q_{i}{k_{t}}^{\top}/\sqrt{d})},\;\;o_{i}=\left(\textstyle\sum_{j=1}^{i}a_{ij}v_{j}\right)W_{o}^{\top}.$ (2)
 Autoregressive LLM Inference.
-The essence of large language model inference is to model the probability distribution $P(x_{1},\ldots,x_{n})$ of a token sequence $x_{1},\ldots,x_{n}$.
-Due to the sequential nature of natural language, prevailing LLMs predominantly employ an autoregressive text generation approach in inference tasks.
-The joint probability of the entire sequence can typically be autoregressively decomposed [5] into a product of conditional probabilities,
-$P(x_{1},\ldots,x_{n})=P(x_{1})P(x_{2}|x_{1})\cdots P(x_{n}|x_{1},\ldots,x_{n-1}).$ (3)
-Specifically, the model generates the output sequence one token at a time.
-When generating each new token, the model considers not only the initial user-provided input but also all previously generated tokens as context.
-For example, when generating the token $x_{i}$, the model utilizes information from tokens $x_{1}$ to $x_{i-1}$.
-While this sequential generation method effectively captures dependencies within the sequence, it introduces significant computational redundancy.
-At each step of generating a new token, attention computations for all preceding tokens (particularly the calculation of $k$ and $v$) must be recalculated.
-This repetitive computation substantially impacts inference efficiency.
+LLM inference models the sequence probability $P(x_{1},\ldots,x_{n})$ through autoregressive decomposition [6].
+The generation process predicts tokens sequentially based on the conditional probability:
+$P(x_{1},\ldots,x_{n})=\prod_{i=1}^{n}P(x_{i}|x_{<i}).$ (3)
+While effective for capturing dependencies, this sequential nature implies that generating $x_{i}$ depends on the full history $x_{<i}$.
+Without caching, the model must re-compute the Key and Value vectors for all preceding tokens at each step.
+This computational redundancy creates a substantial bottleneck for real-time inference.
 KV-cache in LLM Inference.
-In the context of autoregressive generation, the previously generated tokens form the context for predicting the next token.
-To avoid redundant computations, LLMs employ a caching strategy known as the KV-cache.
-For each processed token (either from the input or previously generated by the model), its corresponding $k$ and $v$ vectors, computed by the self-attention layers, are stored in the KV-cache.
-When generating the next token, the model reuses these cached $k$ and $v$ vectors from all preceding tokens, only needing to compute the $q,k,v$ vector for the current prediction step and the $k$ and $v$ vectors for the most recently generated token to add to the cache.
-This significantly accelerates the inference process, especially for long sequences.
+To mitigate the computational redundancy of autoregressive generation, LLMs utilize the KV-cache mechanism [28].
+Instead of recalculating the Key ($K$) and Value ($V$) matrices for the entire context window at each step, the system persists these intermediate states in memory.
+When generating a new token $x_{i}$, the model only computes the current $q_{i},k_{i},v_{i}$, performs attention between $q_{i}$ and the cached history, and subsequently appends the new $(k_{i},v_{i})$ pair to the cache.
+This strategy significantly reduces latency but necessitates managing a persistent state proportional to the sequence length.
 II-B Privacy Attacks against LLMs
-Inversion Attacks.
-While LLMs offer powerful capabilities, their inference process exposes new attack surfaces that can lead to the theft of sensitive user inputs.
-Privacy attacks targeting the inference process can be broadly categorized into two types: (1) Output-based inversion attacks, which aim to reconstruct inputs from the model’s final outputs (e.g., probability distributions or embedding vectors), and (2) Intermediate state-based inversion attacks, which seek to extract information from the model’s internal computational states [15, 21, 34] (e.g., hidden states).
-Early work focused on inverting final representations [15], such as sentence embeddings, with methods like Vec2Text [21].
-However, these approaches overlooked the richer information in deeper layers.
-The Embed Parrot attack [34] demonstrated that even deep hidden states can be exploited to accurately reconstruct user inputs, establishing the entire inference pipeline as a critical attack surface.
-Unlike a hidden state (the fused output of a Transformer layer), the KV-cache contains the more primitive Key (K) and Value (V) vectors that serve as direct inputs to the attention calculation.
-Crucially, the KV-cache is a specialized data structure explicitly designed for persistence and reuse across generation steps to optimize inference.
-This unique role makes it a more potent and exploitable source of privacy leakage than other transient states.
-Side-Channel Attacks Exploiting the KV-cache.
-Currently, research on attacks against the KV-cache has been primarily limited to side-channel attacks.
-A prominent example is PromptPeek [36], which exploits timing discrepancies created by KV-cache sharing in multi-tenant LLM services.
-An attacker observes variations in response latency to determine if their guessed prefix has hit another user’s cached entries, allowing them to reconstruct the prompt without direct data access.
-However, the viability of such timing side-channel attacks is diminished by modern memory management optimizations.
-Contemporary inference engines, particularly those using PagedAttention [14], manage the KV-cache in non-contiguous memory blocks (pages), with a typical default block size of 16 tokens.
-This block-based management means that for a cache hit to occur, an attacker would need to correctly guess an entire block of 16 tokens simultaneously.
-For an LLM with a vocabulary size in the hundreds of thousands, the search space for a single block exceeds $100,000^{16}$, rendering the attack computationally infeasible.
-In contrast, our work considers a different, more direct threat model: if an attacker can gain direct access to the KV-cache content (e.g., through memory access or network eavesdropping), they can bypass these side-channel limitations.
-This direct access enables a more practical collision attack on a per-token basis, reducing the attack complexity to a feasible level and exposing a more fundamental vulnerability.
-Given that the KV-cache is commonly stored and transmitted in plaintext to preserve performance, this paper aims to systematically explore this underexplored problem of direct, content-based attacks.
-III Attack Landscape: Inferring Private User Inputs from KV-cache
-While KV-cache is crucial for enhancing the inference efficiency of LLMs, it concurrently introduces novel privacy vulnerabilities.
-The contextual information stored within KV-cache can contain private user data.
-Specifically, entries in the KV-cache often correspond directly to tokens in the user’s input sequence.
-An attacker gaining access to this KV-cache information could potentially infer the original user input, thereby compromising user privacy.
-In this section, we aim to study an important research question: is an attacker able to conduct a prompt reconstruction attack to recover the user input from the KV-cache?
-We first present the threat model of the attack.
-Subsequently, we propose two distinct attacks to achieve this goal.
+II-B1 Inference-Phase Inversion Attacks
+Inference-time privacy attacks are categorized into: (1) Output-based, reconstructing inputs from final probabilities or embeddings [19, 25]; and (2) Intermediate-state-based, exploiting internal representations [37].
+While early works like Vec2Text [25] focused on sentence embeddings, Embed Parrot [37] demonstrated that deep hidden states allow for more accurate reconstruction.
+However, current research often overlooks the KV-cache.
+Unlike fused hidden states, the KV-cache retains the primitive Key ($K$) and Value ($V$) vectors—the raw inputs to the attention mechanism.
+Crucially, to support autoregressive generation, the KV-cache is architecturally designed for persistence, making it a richer and more exploitable surface than transient computational states.
+II-B2 Limitations of Side-Channel Attacks
+Existing threats to the KV-cache are primarily side-channel attacks, such as PromptPeek [39], which infers prompts by measuring latency variations caused by cache sharing.
+However, the practicality of such timing attacks is severely mitigated by modern memory management.
+Systems like PagedAttention [17] manage the KV-cache in non-contiguous blocks (defaulting to 16 tokens).
+To trigger a cache hit (and thus a timing signal), an adversary must correctly guess an entire block simultaneously.
+With a vocabulary size $|V|\approx 10^{5}$, the search space explodes to $|V|^{16}$, rendering brute-force side-channels computationally infeasible.
+In contrast, we address the direct access threat model.
+Given that KV-caches are often processed in plaintext for performance, a compromised system allows adversaries to bypass probabilistic guessing and perform precise, per-token collision attacks.
+III Attack Landscape: Input Reconstruction from KV-cache
+While the KV-cache serves as a cornerstone for efficient LLM inference, it concurrently introduces a critical and underexplored privacy surface.
+Unlike deep hidden states that represent fused semantic information, entries in the KV-cache maintain a direct, element-wise correspondence to the tokens in the user’s input sequence.
+This structural characteristic implies that an adversary gaining access to the KV-cache does not merely obtain abstract embeddings, but potentially holds the raw materials to reverse-engineer the original user prompt.
+In this section, we investigate the central research question: Is an adversary able to reconstruct user inputs from the KV-cache?
+We first define a realistic threat model grounded in modern cloud architectures, and subsequently detail three distinct attack vectors designed to exploit this vulnerabilities.
 III-A Threat Model
-In our threat model, we assume that an LLM server has deployed an open-source LLM based service.
-After a user uses the service with private user inputs, it generates a large number of KV-caches.
-The attacker gains access to the KV-caches and tries to infer the private inputs, as showed in Figure 2.
-Figure 2: The attacker restores user input based on model parameters and leaked KV-cache.
-Attacker’s Objective.
-The primary objective of the attacker is to recover the user input prompt from the accessed KV-cache.
-In this paper, we assume that the attacker attempts to restore the user input verbatim, which may significantly compromise the privacy of the user.
-Attacker’s Capabilities. We consider an attacker with relatively strong yet still practical capabilities, as follows.
-  • Access to KV-cache.
-The attacker can obtain the KV-caches stored by an LLM server.
-This can be achieved by eavesdropping on the communication channel between nodes (e.g.
-through the LeftoverLocals attack [28]), reading persisted KV-cache from storage, or other unspecified means of accessing the KV-cache pool.
-  • Access to the foundation model on which the deployed LLM is developed.
-We assume a gray-box setting where the attacker can gain access to the foundation model’s weights on which the deployed LLM is developed [^1] This assumption is realistic given that now most LLM services are developed on open-source foundation models.
-First, disclosing the underlying foundation model is often a fundamental requirement of open-source licenses.
-Second, even if model information isn’t publicly disclosed, an attacker could identify the specific foundation model via model fingerprinting attacks [23, 7]..
-The deployed model could be a fine-tuned version of the foundation model.
+This work focuses on the threat landscape of LLM inference services within a confidential computing paradigm. We consider the inference service provider as the adversary.
+Adversary’s Objective.
+The primary objective of the adversary is to recover the user input prompt from the accessed KV-cache.
+We target verbatim reconstruction because it represents the most severe breach of privacy, allowing the extraction of exact credentials, PII, or proprietary logic contained in the input.
+Adversary’s Capabilities.
+We assume the adversary can obtain both the KV-cache and the model weights, but does not observe the ephemeral runtime activations within the GPU registers.
+These assumptions are well-grounded in the realities of current confidential computing inference paradigms:
+  • KV-cache Access.
+The adversary is capable of obtaining the KV-cache stored by the LLM server.
+This is a realistic assumption, as high-throughput and scalable cloud-native services deliberately externalize the large-scale KV-cache to non-secure memory or persistent storage pools to meet performance demands [43].
+We thus consider this access a result of an intentional performance-security trade-off, not a traditional vulnerability.
+  • Foundation Model Access.
+We assume a gray-box setting where the adversary can access the foundation model weights.
+This access is realistic through two primary paths: (1) For closed-source services, the CSP inherently owns the model and possesses the weights.
+(2) For services built on open-source models, the provider may be required to disclose the base model for licensing compliance, or the model can be identified using fingerprinting techniques [27, 8].
 III-B Input Reconstruction Attacks from KV-cache
+(a) KV-cache Inversion Attack.
+(b) KV-cache Collision Attack.
+(c) KV-cache Injection Attack.
+Figure 2: Workflow of the three proposed KV-cache input reconstruction attacks.
 This section investigates the risk of user input leakage from KV-cache data during LLM inference.
@@
 These attacks differ in complexity, applicability, and exploited vulnerabilities, collectively illustrating the threat landscape.
-(a) KV-cache Inversion Attack.
-(b) KV-cache Collision Attack.
-(c) KV-cache Injection Attack.
-Figure 3: Workflow of input reconstruction attacks.
 III-B1 KV-cache Inversion Attack
-The KV-cache Inversion Attack, as showed in Figure 3a, is a direct method that attempts to reverse-engineer the input to the attention layer by leveraging the KV-cache data in conjunction with the model’s known attention projection matrices.
-As defined in the forward computation (Eq.
-(1)), an attacker who obtains the KV-cache and knows the model’s attention layer parameters $W_{k}$ and $W_{v}$ could theoretically perform a reverse calculation to recover the attention layer’s input $x_{i}$, as Eq.
-(4).
-$x_{i}=k_{i}({R}_{\Theta,i}^{d})^{-1}(W_{k}^{\top})^{-1},\quad x_{i}=v_{i}(W_{v}^{\top})^{-1}.$ (4)
-However, the practicality of this direct inversion attack is severely constrained by two fundamental factors.
-First, it requires the attention projection matrices (e.g., $W_{k},W_{v}$) to be invertible.
-This condition is met by some earlier models using standard Multi-Head Attention (MHA), such as LLaMA-7B [32], but not by many contemporary models.
-Architectures like LLaMA 3 [9], Qwen [37], and DeepSeek [20] employ optimizations like Grouped-Query Attention (GQA)[3] or Multi-Head Latent Attention (MLA)[26], which often result in non-square, non-invertible projection matrices, rendering the attack impractical.
-Second, the attack’s effectiveness is largely confined to the first decoder layer’s KV-cache, as its inversion directly yields the input sequence’s embeddings.
-In contrast, inverting from deeper layers only recovers intermediate hidden states, from which reconstructing the original text is a significantly more complex and data-intensive problem often requiring auxiliary models [34].
-Due to these significant limitations, the KV-cache Inversion Attack, while theoretically straightforward, has a narrow scope of applicability and does not pose a threat to most modern LLM architectures.
+As illustrated in Figure 2a, the KV-cache Inversion Attack functions as a baseline algebraic adversary.
+It attempts to mathematically invert the attention mechanism by exploiting the linear projection of Key ($k$) and Value ($v$) vectors.
+Attack Formulation.
+Based on the forward pass definition in Eq.
+(1), an adversary possessing the plaintext KV-cache and the model weights ($W_{k},W_{v}$) can theoretically recover the input state $x_{i}$ of the attention layer via matrix inversion:
+$x_{i}=k_{i}({R}_{\Theta,i}^{d})^{-1}(W_{k}^{\top})^{-1},\;\;x_{i}=v_{i}(W_{v}^{\top})^{-1}.$ (4)
+Architectural and Semantic Constraints.
+While algebraically sound, the feasibility of this attack in production environments is severely limited by two fundamental constraints: (1) The attack strictly requires the projection matrices $W_{k}$ and $W_{v}$ to be square and full-rank invertible.
+This condition holds for legacy Multi-Head Attention (MHA) [36] architectures (e.g., LLaMA-7B).
+However, modern State-of-the-Art (SOTA) models, including LLaMA-3 [12], Qwen [40], and DeepSeek [24], adopt efficiency optimizations such as Grouped-Query Attention (GQA) [3] or Multi-Head Latent Attention (MLA) [30].
+These mechanisms typically project inputs into lower-dimensional subspaces, resulting in non-square matrices where unique inversion is mathematically impossible (i.e., the problem becomes ill-posed).
+(2) the attack’s effectiveness is largely confined to the first decoder layer’s KV-cache, as its inversion directly yields the input sequence’s embeddings, allowing for precise mapping back to the vocabulary.
+In contrast, applying Eq.
+(4) to deeper layers recovers intermediate hidden states.
+These states represent semantically fused contextual information rather than discrete tokens.
+Reversing these deep representations into text is a non-trivial problem that necessitates training auxiliary inversion models [37], thereby increasing the adversary’s computational cost and reducing fidelity.
+Consequently, while the Inversion Attack serves as a theoretical proof-of-concept for KV-cache leakage, its reliance on specific architectural properties (MHA) and layer positions (Layer 0) limits its utility against modern, deep LLMs.
+This motivates the need for the more robust attacks.
 III-B2 KV-cache Collision Attack
-Unlike the KV-cache Inversion Attack, the KV-cache Collision Attack operates by iteratively generating KV-cache entries for candidate tokens from the model’s vocabulary.
-For each candidate token appended to a partially reconstructed input sequence, the attack performs local model inference, generates the corresponding KV-cache, and compares it against the intercepted KV-cache to determine if the candidate token is correct for the current position.
-Fundamental Principle of the Collision Attack.
-The collision attack reconstructs user input by identifying the token at each position that minimizes a distance metric (e.g., L2 distance) between a locally generated KV-cache and the leaked KV-cache.
-Its effectiveness hinges on the assumption that the distance for the correct token, $dis_{target}$, is statistically separable from the distances of all other incorrect tokens from the vocabulary, $dis_{other}$, allowing for its unambiguous identification.
-General Attack Procedure.
-As illustrated in Figure 3b, the attack reconstructs the user’s input token-by-token.
-After initially intercepting the target KV-cache from the cloud service, the attacker enters an iterative loop for each unknown token position.
-First, guided by the already-reconstructed prefix, the attacker’s local model instance predicts a probabilistically-ranked list of candidate tokens.
-Next, for each high-probability candidate, the attacker performs a forward pass to generate a local KV-cache segment and compares it against the corresponding segment of the target KV-cache to test for a match (a “collision”).
-Upon a successful match, the correct token is confirmed and appended to the reconstructed prefix, and the attacker proceeds to the next position.
-This process continues until the entire input sequence is recovered.
-Implementation Optimizations for the Collision Attack.
-A naive collision attack that processes the entire model vocabulary is impractical due to prohibitive VRAM consumption and execution time.
-We therefore introduce some synergistic optimizations to make the attack practical.
-First, instead of analyzing all prioritized candidates, we apply batched statistical validation, testing a subset of candidate tokens and computing the distribution of their distances ($dis_{batch}$) from the leaked KV-cache.
-The correct token is then identified as any distinct statistical outlier within this batch, under the assumption that the batch provides a representative sample of the distance distribution for incorrect tokens.
-Second, we employ a probability-guided search.
-Using the model’s own predictions based on the already reconstructed prefix, we prioritize collision tests for high-likelihood candidate tokens.
-Furthermore, we prune the search space by truncating the vocabulary and discarding tokens with very low probabilities.
-Together, these techniques reduce the memory footprint and bring the attack latency down to the order of seconds.
-Enhancing Collision Attack with Prior Knowledge.
-A key challenge in the collision attack is setting an optimal classification threshold to distinguish the correct token’s distance ($dis_{target}$) from the distribution of incorrect token distances ($dis_{other}$).
-A basic attacker, lacking ground-truth data, must rely on a static, heuristic threshold, typically derived from the statistics of observed incorrect token distances (e.g., $\mu_{other}-3\sigma_{other}$).
-However, an adversary capable of performing chosen-plaintext attacks can generate KV-caches for known inputs.
-This prior knowledge enables a more precise statistical characterization of both the $dis_{target}$ and $dis_{other}$ distributions, allowing for a more effective, adaptive threshold that significantly boosts reconstruction accuracy.
-The need for an adaptive threshold is formalized by the success probability for a token at rank $r$ in the probability-sorted vocabulary, as defined in Eq. (5):
-$P(\text{success}|t)=P(dis_{other}>t)^{r-1}\times P(dis_{target}<t).$ (5)
-This equation demonstrates that the optimal threshold $t$ is dependent on the token’s expected rank $r$, as it must balance the trade-off between correctly rejecting the r1r-1 preceding (incorrect) candidates and accepting the true token.
-A static heuristic, by contrast, cannot effectively manage this rank-dependent trade-off.
-The KV-cache Collision Attack overcomes the limitations of the Inversion Attack concerning specific KV-cache and model attention mechanism requirements.
-An attacker who intercepts KV-cache from potentially any layer of various LLM architectures can, in principle, employ this collision-based approach to reconstruct user inputs.
-The practical effectiveness and performance of this attack will be demonstrated and evaluated in subsequent experimental sections.
+Unlike the Inversion Attack which suffers from architectural constraints, the KV-cache Collision Attack is a universal, forward-matching adversary applicable to any layer.
+Instead of algebraically reversing the projection, this attack reframes input reconstruction as a search problem: identifying the token in the vocabulary that, when processed by a local model, yields a KV-cache entry most similar to the intercepted target.
+Fundamental Principle: Reconstruction as Optimization.
+The core premise is that the KV-cache exhibits distance-preserving properties for identical tokens under the same context.
+We formalize the attack as finding the optimal token $t^{*}$ at position $i$ that minimizes the distance metric $\mathcal{D}$ (e.g., Frobenius norm) between the leaked cache $K_{\text{leaked}}$ and the local generation $K_{\text{local}}(t)$:
+$t^{*}_{i}=\mathop{\arg\min}_{t\in\mathcal{V}}\mathcal{D}\left(K_{\text{leaked}}^{(i)},K_{\text{local}}^{(i)}(t|x_{<i})\right).$ (5)
+The attack’s success hinges on statistical separability: the distance for the ground-truth token, $d_{\text{target}}$, must be statistically distinguishable from the distribution of distances for incorrect tokens, $d_{\text{other}}$.
+Attack Procedure.
+As illustrated in Figure 2b, the attack proceeds iteratively token-by-token.
+For each unknown position $i$: (1) Candidate Generation: The adversary selects candidate tokens from the model’s vocabulary $\mathcal{V}$ to test.
+(2) Local Simulation: For each candidate token $t$, the adversary performs a forward pass using their local model instance to generate the corresponding KV-cache entry $K_{\text{local}}^{(i)}(t)$.
+(3) Collision Matching: The adversary computes the distance $\mathcal{D}$ between the local entry and the intercepted target $K_{\text{leaked}}^{(i)}$.
+The candidate that minimizes this distance (i.e., creates a “collision” with the leaked state) is identified as the correct token $x_{i}$, appended to the sequence, and the process advances to $i+1$.
+Implementation Optimizations.
+A naive exhaustive search over the entire vocabulary ($\mathcal{V}\approx 10^{5}$) entails prohibitive latency and VRAM consumption, as verifying the global minimum distance requires inferring every candidate token.
+To render the attack practical, we implement three synergistic optimizations that transform the problem from a global search into an efficient sequential decision process: (1) Batched Outlier Detection (Enabling Early Exit).
+Instead of calculating distances for the full vocabulary to find the minimum, we process candidates in batches.
+We fundamentally shift the decision logic: rather than comparing a candidate against all other tokens in $\mathcal{V}$, we determine if a candidate constitutes a statistical outlier within its current batch (i.e., $d_{t}<\mu_{\text{batch}}-3\sigma_{\text{batch}}$).
+Crucially, this allows for an early exit: once a collision is statistically confirmed within a batch, the search terminates immediately without evaluating the remaining vocabulary.
+(2) Probability-Guided Prioritization.
+To maximize the efficacy of the early exit mechanism, we employ a probability-guided search.
+We sort the candidate tokens based on the model’s predicted probabilities $P(x_{i}|x_{<i})$.
+This reordering ensures that the correct token—which typically carries a high probability—is placed in the first few batches.
+Consequently, the collision is triggered near the start of the search process, drastically reducing the average number of inferences required.
+(3) Search Space Pruning.
+Distinct from prioritization, we explicitly truncate the search space by discarding the “long tail” of tokens with negligible probabilities (e.g., keeping only the top-$k\%$ candidates).
+While prioritization accelerates the average case, pruning bounds the worst-case complexity.
+It prevents the adversary from wasting resources iterating through tens of thousands of implausible tokens in the rare event that the target is not found in the high-probability regions.
+Enhancement via Chosen-Plaintext Attack (CPA).
+A critical challenge is determining the optimal decision threshold $\tau$.
+A static heuristic (e.g., $3\sigma$-rule) is often suboptimal due to the varying entropy of different tokens.
+However, an adversary with CPA capabilities—who can feed known inputs to the model and observe the resulting KV-cache—can profile the precise statistical behavior of $d_{\text{target}}$ and $d_{\text{other}}$.
+This prior knowledge allows for an adaptive thresholding strategy. We formulate the success probability for a candidate at rank $r$ as:
+$P(\text{success}|r)=P(d_{\text{other}}>\tau)^{r-1}\times P(d_{\text{target}}<\tau).$ (6)
+This formulation explicitly captures the trade-off: a stricter threshold minimizes False Positives (accepting an incorrect predecessor) but risks False Negatives (rejecting the target).
+By dynamically adjusting $\tau$ based on the expected rank $r$ and the profiled distributions from CPA, the adversary maximizes $P(\text{success})$, significantly boosting reconstruction fidelity compared to static baselines.
 III-B3 KV-cache Injection Attack
-The KV-cache Injection Attack leverages the LLM’s own semantic capabilities to exfiltrate information, operating differently from direct state-matching methods.
-The KV-cache contains substantial information from user-model interactions, however, itself (comprising only $k$ and $v$ vectors for past tokens) lacks the $q$ vectors for the current generation step and thus cannot be directly used to continue model inference without a new input token.
-The core principle involves an attacker appending a crafted instruction (e.g., “Repeat the previous content.”) to an intercepted KV-cache, as showed in Figure 3c.
-This action provides the necessary query, key, and value vectors to resume inference from the compromised state, effectively tricking a standard LLM into executing the command within the context of the user’s private data.
-This can compel the model to “echo” or regenerate the sensitive input stored implicitly within the cache.
-While this method is generally less precise than the collision attack, its primary advantage is speed, requiring only a single forward inference pass.
-The viability of this attack vector reveals a crucial requirement for any comprehensive defense: a protected KV-cache must be rendered semantically unintelligible to prevent an unauthorized LLM from meaningfully parsing its contents or continuing generation from it.
-We conduct comprehensive evaluations on these three attacks in Section V.
-The experimental results demonstrate the practicality and effectiveness of input reconstruction attacks from KV-cache, underscoring the urgent need for developing an effective defense technique.
-IV KV-Cloak: A Defense Mechanism for
-KV-cache Obfuscation
-IV-A Motivation for KV-Cloak
-Limitations of Existing Privacy-Preserving Techniques.
-While several techniques exist for data privacy, they prove impractical or inadequate for securing the LLM KV-cache, highlighting the need for a specialized solution.
-The introduction to existing privacy-preserving techniques and their limitations is as follows.
-  • Cryptographic Methods: Standard cryptographic techniques [4, 31], such as symmetric AES encryption, provide confidentiality by ensuring data remains in ciphertext during storage and transmission, requiring decryption only for legitimate use.
-A more advanced approach, Homomorphic Encryption (HE) [22, 2], allows for direct computation on encrypted data, theoretically enabling parts of the attention mechanism to operate on the KV-cache without ever decrypting it.
-However, despite their strong security guarantees, the immense computational overhead and latency introduced by these methods are prohibitive.
-Given that the KV-cache can be tens or hundreds of gigabytes, applying full encryption or HE is unsuitable for the high-throughput, low-latency requirements of LLM inference.
-  • Differential Privacy (DP): Differential Privacy is a rigorous model that protects individual privacy by injecting carefully calibrated random noise into data or query results [1, 16, 39].
-The fundamental goal is to make the output of a computation statistically insensitive to the presence or absence of any single individual’s data in the input dataset.
-In the context of the KV-cache, this would involve adding noise to the Key and Value vectors.
-The primary limitation of this approach is the inherent trade-off between privacy and utility.
-To achieve a meaningful level of privacy, the required amount of noise would significantly degrade the LLM’s inference accuracy to an unacceptable degree.
-  • KV-Shield: KV-Shield [38] is a lightweight obfuscation scheme and, to our best knowledge, the only privacy-preserving method designed specifically for the KV-cache.
-Its process involves synchronously permuting the rows of the model’s attention layer parameters ($W_{q},W_{k},W_{v}$).
-This operation effectively shuffles the element positions within the resulting KV-cache vectors, obfuscating them before an attention score is calculated.
-The obfuscated attention output is then de-obfuscated for subsequent steps.
-However, this approach has critical flaws.
-From a security perspective, its fixed shuffling pattern does not alter the statistical properties of the data, leaving it vulnerable to collision attacks.
-For compatibility, its technique is incompatible with modern LLM architectures that utilize features like Rotary Positional Embedding (RoPE), as the shuffling interferes with the position-dependent calculations.
-Design Objectives.
-The shortcomings of existing methods reveal an urgent need for a defense mechanism that is both secure and practical.
-To be effective, such a solution must achieve three key objectives:
-  • High Security: It must provide robust protection against targeted reconstruction attacks.
-  • Minimal Accuracy Impact: It must operate without causing any degradation in the model’s output accuracy.
-  • Low Performance Overhead: It must introduce only negligible latency into the inference pipeline.
-To meet these challenges, we propose KV-Cloak, a novel defense mechanism designed to provide strong, lossless, and efficient protection for the KV-cache.
-IV-B A Naive Defense: Obfuscation via Reversible Linear Transforms
-We first evaluate a baseline defense that employs a reversible linear transform to obscure the statistical properties of the $k$ and $v$ vectors.
-In the context of modern inference frameworks like vLLM, we represent all key vectors for a single attention head within a KV-cache block as a matrix $K\in\mathbb{R}^{b\times d}$, where $b$ is the number of vectors per block and $d$ is the attention head dimension.
-As the obfuscation method is analogous for both keys and values, we use $K$ as our canonical example.
-The transformation is defined as:
-$K^{\prime}=SKM,$ (6)
-where $S\in\mathbb{R}^{b\times b}$ and $M\in\mathbb{R}^{d\times d}$ are secret, randomly generated, and invertible matrices.
-The objective is to transform the original matrix $K$ without altering its dimensions, preventing an adversary from directly recovering its contents.
-Vulnerability Analysis.
-Despite its simplicity, this linear transform is fundamentally insecure under a Chosen-Plaintext Attack (CPA).
-While an adversary cannot craft an arbitrary matrix $K$ from scratch—as its contents are determined by the model’s embedding and projection layers—they can still exert significant influence through carefully constructed inputs.
-This capability is sufficient to break the scheme.
-The fixed linear nature of the transformation makes it highly susceptible to algebraic cryptanalysis.
-An adversary can mount a differential attack by choosing two plaintexts, $K_{1}$ and $K_{2}$, thereby controlling the difference $\Delta K=K_{1}-K_{2}$ and observing the corresponding output difference $\Delta K^{\prime}=S(\Delta K)M$.
-By systematically choosing $\Delta K$ to be a series of standard basis matrices (i.e., matrices with a single non-zero entry), the adversary can isolate the columns of $S$ and the rows of $M$.
-This process allows for the full recovery of the secret matrices $S$ and $M$ (up to a trivial scaling ambiguity), leading to a complete compromise of the scheme.
-IV-C The Improved Defense: One-Time Pad Block-wise Shuffling
-Eliminating Redundant Positional Information in KV-cache.
-To enhance the security of the obfuscation scheme, we introduce additional randomness by eliminating redundant positional information inherent to KV-cache storage.
-During attention mechanism computations, the physical ordering of $k,v$ vectors in the KV-cache is an unnecessary redundancy for inference.
-In reality, positional information critical for computation is already embedded into the elements of $q$ and $k$ vectors through rotary positional encoding.
-Figure 4: Shuffling $k,v$ vectors within a KV-cache block preserves attention computation results while disrupting positional order.
-By randomly permuting $k,v$ within each block while maintaining their correspondence to input tokens $x$, we eliminate positional clues that attackers could exploit, and the results of the attention mechanism were not affected (as Figure 4).
-This permutation increases the computational complexity for attackers to match plaintext-ciphertext pairs by a factor of $b!$, where $b$ denotes the block size (default $b=16$ in vLLM).
-The enhanced obfuscation process employs a one-time pad permutation matrix $\hat{P}$ to reorder vectors within blocks before applying invertible matrices $S$ and $M$:
-$K^{\prime}=S\hat{P}KM.$ (7)
-Crucially, $\hat{P}$ does not need to be stored as a key since the deobfuscated $\hat{P}K$ can directly participate in subsequent computations without reconstruction.
-This design choice preserves computational efficiency.
-Additive Mask Matrices with Position-Identifiable Features.
-However, when the original $K$ matrix has rank $1$ (i.e., all row vectors are identical), the entropy of the permutation $\hat{P}$ will collapse.
-To prevent rank deficiency, KV-Cloak introduce an additive masking matrix $A$ before permutation.
-Thereby, we can ensure the transformed matrix $(K+A)$ maintains sufficient rank to preserve cryptographic strength of $\hat{P}$.
-Since the one-time pad permutation matrix $\hat{P}$ is not stored, computing $\hat{P}A$ during inverse obfuscation becomes infeasible, and thus $\hat{P}K=\hat{P}(K+A)-\hat{P}A$ cannot be derived from $\hat{P}(K+A)$.
-To avoid storing the one-time pad permutation matrix $\hat{P}$, we exploit the sparsity and small-elements properties of $k$ and $v$ vectors to design the additive matrix $A$ with additional structural constraints.
-This enables us to infer $\hat{P}A$ based on the data characteristics of $\hat{P}(K+A)$.
-Figure 5: Using an excerpt from “The Bitter Lesson” [30] (see Appendix -B) as the input of LLaMA-3.2-1b, the KV-cache data element distribution generated during the inference process.
-As shown in Figure 5, we observe that the elements of $k$ and $v$ vectors typically remain below a small threshold $\theta_{K}$ (e.g., 100).
-This allows us to construct a matrix $A$ where positional information is embedded into $K+A$ via additive operations using values significantly larger than $\theta_{K}$ in specific row vectors.
-After applying row permutations, the positional identifiers in $\hat{P}(K+A)$ remain identifiable.
-During inverse obfuscation, the positional information of $\hat{P}A$ can be directly identified from $\hat{P}(K+A)$ by detecting these outliers.
-Subtracting $\hat{P}A$ from $\hat{P}(K+A)$ yields $\hat{P}K$ for subsequent inference, drastically reducing storage overhead for the one-time pad key.
-The complete obfuscation transformation, as shown in Eq. (8), incorporates additional matrices $\hat{P}$ and $A$ alongside invertible linear transformations $S$ and $M$:
-$K^{\prime}=S\hat{P}(K+A)M,$ (8)
-[…]
-IV-D Implicit Obfuscation via Operator Fusion
-To mitigate the performance overhead introduced by the obfuscation and de-obfuscation processes during runtime, we now demonstrate how the obfuscation operators can be fused directly into the model’s weight matrices.
-This allows the primary cost of obfuscation to be paid offline, resulting in a minimal impact on online inference latency.
-Fusing the Obfuscation Operator into Model Weights.
-We define three transformed vectors $(q^{m},k^{m},v^{m})$ and a transformed output projection matrix $W_{o}^{m}$ using two secret, invertible matrices, $M_{1}$ and $M_{2}$.
-Let $q,k,v$ be the original query, key, and value vectors, and let $W_{o}$ be the original output projection matrix.
-The transformations are defined as follows:
-$\left\{\begin{array}[]{l}q^{m}=q(M_{1}^{-1})^{\top},\\ k^{m}=kM_{1},\\ v^{m}=vM_{2},\\ W_{o}^{m}=W_{o}(M_{2}^{-1})^{\top}.\end{array}\right$. (9)
-These transformations are specifically designed to leave the core attention mechanism computations invariant. The attention scores remain unchanged, as shown by the dot product:
-$q_{i}^{m}(k_{j}^{m})^{\top}$ $=\left(q_{i}(M_{1}^{-1})^{\top}\right)\left(k_{j}M_{1}\right)^{\top}=q_{i}(M_{1}^{-1})^{\top}M_{1}^{\top}k_{j}^{\top}$ (10)
-$=q_{i}(M_{1}M_{1}^{-1})^{\top}k_{j}^{\top}=q_{i}k_{j}^{\top}.$
-Similarly, the output of the attention head after projection remains identical:
-$v_{j}^{m}(W_{o}^{m})^{\top}$ $=(v_{j}M_{2})\left(W_{o}(M_{2}^{-1})^{\top}\right)^{\top}$ (11)
-$=v_{j}M_{2}(M_{2}^{-1})W_{o}^{\top}=v_{j}W_{o}^{\top}.$
-This invariance guarantees that the model’s output is mathematically identical to that of the original, unprotected model, ensuring lossless accuracy.
-Next, we incorporate the projection from the input embedding $x$ and the application of RoPE, denoted by the position-dependent matrix $R_{\Theta,i}^{d}$.
-The equations for the transformed vectors become:
-$\left\{\begin{array}[]{l}q^{m}=xW_{q}^{\top}R_{\Theta,i}^{d}(M_{1}^{-1})^{\top}\\ k^{m}=xW_{k}^{\top}R_{\Theta,i}^{d}M_{1}\\ v^{m}=xW_{v}^{\top}M_{2}\\ W_{o}^{m}=W_{o}(M_{2}^{-1})^{\top}\end{array}\right.$.
-(12)
-From these formulations, we can observe the path to operator fusion.
-The secret matrices $M_{2}$ and $M_{2}^{-1}$, which obfuscate the value vectors, can be straightforwardly fused into the model’s weights by pre-computing new weight matrices $(W_{v}^{m})^{\top}=W_{v}^{\top}M_{2}$ and $W_{o}^{m}=W_{o}(M_{2}^{-1})^{\top}$.
-However, fusing $M_{1}$ and $M_{1}^{-1}$ into $W_{k}$ and $W_{q}$ is more complex due to the intermediate application of the position-dependent RoPE matrix $R_{\Theta,i}^{d}$.
-The fusion becomes feasible under a specific algebraic condition: if the secret matrix $M_{1}$ and the RoPE matrix $R_{\Theta,i}^{d}$ commute, i.e., $R_{\Theta,i}^{d}M_{1}=M_{1}R_{\Theta,i}^{d}$.
-If this condition holds, we can rewrite the expression for $k^{m}$:
-$k^{m}=xW_{k}^{\top}R_{\Theta,i}^{d}M_{1}=x(W_{k}^{\top}M_{1})R_{\Theta,i}^{d}.$ (13)
-This allows us to pre-compute a new, position-independent weight matrix $(W_{k}^{m})^{\top}=W_{k}^{\top}M_{1}$. A similar argument applies to fusing $(M_{1}^{-1})^{\top}$ into $W_{q}$.
-Therefore, the viability of this operator fusion approach hinges on our ability to construct a secret matrix $M_{1}$ that commutes with the RoPE matrices.
-In Appendix -A, we formally prove that this commutativity is achieved when $M_{1}$ is structured as a block-diagonal matrix composed of $2\times 2$ rotation-scaling sub-matrices, a structure analogous to that of RoPE itself.
-Furthermore, to ensure the reversibility of the entire transformation for lossless de-obfuscation, $M_{1}$ must be constructed to be invertible.
-By constraining $M_{2}$ to also be a random, invertible rotation-scaling matrix, and by carefully controlling the range of rotation-scaling coefficients in both $M_{1}$ and $M_{2}$, we ensure that the transformed padding tokens (which involve the affine transformation with matrix $A$) remain identifiable as outliers.
-These conditions allow the obfuscation operators ($M_{1},M_{2}$) and their corresponding de-obfuscation counterparts ($M_{1}^{-1},M_{2}^{-1}$) to be fused into the attention layer parameters $W_{q},W_{k},W_{v},\text{and }W_{o}$.
-Let the model parameters after operator fusion be denoted by $W_{q}^{m},W_{k}^{m},W_{v}^{m},W_{o}^{m}$.
-The new, fused weights are pre-computed as follows:
-$\left\{\begin{aligned} W_{q}^{m}&=M_{1}^{-1}W_{q}\\ W_{k}^{m}&=M_{1}^{\top}W_{k}\\ W_{v}^{m}&=M_{2}^{\top}W_{v}\\ W_{o}^{m}&=W_{o}(M_{2}^{-1})^{\top}\end{aligned}\right.$. (14)
-By deploying the LLM with these pre-computed fused weights, we effectively integrate the random invertible block-diagonal matrix $M_{1}$ and the general random invertible matrix $M_{2}\in\mathbb{R}^{d\times d}$ into the attention layer parameters.
-This eliminates the need for explicit online matrix multiplications with $M_{1}$ and $M_{2}$, significantly reducing performance overhead.
-Ultimately, the KV-Cloak obfuscation formula, leveraging operator fusion, is applied as follows:
-$K^{\prime}=S\hat{P}(K^{m}+A),$ (15)
-where $K^{m}=KM$ represents the KV-cache generated by the inference process using the fused model weights.
-It is crucial to control the range of the rotation and scaling factors within the matrix $M$ to ensure that the additive mask $A$ can still reliably embed its positional information via the outlier mechanism.
-TABLE I: Effectiveness of the inversion, collision, and injection attacks against the unprotected KV-cache from different model layers.
-Here, “collision+” denotes the collision attack enhanced with prior knowledge.
-Model Metric Inversion Collision Collision+ Injection
+Unlike Inversion and Collision attacks which rely on precise algebraic state matching, the KV-cache Injection Attack exploits the semantic instruction-following capabilities of LLMs to exfiltrate private contexts.
+Architectural Constraint: The Missing Query.
+Architecturally, a stand-alone KV-cache is computationally dormant.
+The self-attention mechanism requires a current Query vector ($Q$) to drive the attention scoring ($\text{Softmax}(QK^{\top})$) and state transitions.
+Consequently, an adversary cannot simply “resume” inference from a stolen cache; a new input stimulus is required to generate the necessary activation state.
+Attack Mechanism: Contextual Hijacking.
+To overcome this, we weaponize the LLM’s instruction-following nature.
+As shown in Figure 2c, the adversary appends a crafted directive (e.g., “Repeat the previous content.”) to the intercepted KV-cache.
+This injection serves two purposes: (1) Stimulus Generation.
+The directive creates the requisite $Q$ vectors to activate the attention mechanism.
+(2) Forced Exfiltration.
+The LLM attends to the stolen $K/V$ pairs as historical context.
+Bound by its alignment training, the model executes the instruction, effectively “echoing” or summarizing the latent private information stored in the cache.
+Strategic Advantages: Robustness and Efficiency.
+This attack offers distinct advantages in robustness compared to the Collision Attack.
+Specifically, it remains effective against KV-cache eviction strategies like H2O [46].
+While such lossy compression breaks the strict mathematical correspondence required for algebraic attacks, it typically preserves the semantic gist of the context.
+The Injection Attack successfully exploits these residual semantic vectors to hallucinate or reconstruct private data.
+Furthermore, the attack is highly efficient, requiring only a single generation pass.
+This vector underscores that a comprehensive defense must render the KV-cache semantically unintelligible, not just algebraically obfuscated.
+IV Evaluation of Attacks
+IV-A Experimental Setup
+Models.
+To demonstrate attack universality, we select seven state-of-the-art LLMs spanning diverse parameter scales (1B–8B) and attention mechanisms.
+Our evaluation prioritizes modern GQA architectures, including the LLaMA-3.2 series (1B & 3B-Instruct), LLaMA-3.1-8B [12], and Qwen2.5-Math-7B [40].
+Crucially, we incorporate DeepSeek-R1-Distill-LLaMA-8B [13] (LLaMA-3.1-8B-Distilled) to assess robustness against fine-tuned models where weight parameters diverge from their base counterparts.
+Finally, to verify generalizability across architectural evolutions, we evaluate legacy MHA models, specifically LLaMA-7B and LLaMA-2-7B.
+Datasets.
+To simulate realistic privacy leakage scenarios, we utilize the LMSYS-Chat-1M dataset [48], which comprises real-world user interactions collected from inference services.
+We construct a test set by randomly sampling 1,000 instances, ensuring a rich variety of dialogue and instruction-following contexts.
+Analysis of attack generalization across domain-specific datasets is provided in Appendix -C3.
+Evaluation Metrics.
+BERTScore [45] and ROUGE-L [23] are used to measure the similarity between the original input and the text reconstructed from the KV-cache.
+BERTScore, based on the all-mpnet-base-v2 model, is better at capturing semantic similarity, while ROUGE-L primarily reflects lexical-level precision and recall.
+IV-B Attack Effectiveness
+TABLE I: Reconstruction fidelity (BERTScore and ROUGE-L) of attacks against unprotected KV-cache across different models and layers. “Collision+” denotes the attack enhanced with prior knowledge.
+Type Model Metric Inversion Collision Collision+ Injection
 First Mid Last First Mid Last First Mid Last All
-LLaMA-7B BERTScore ($\uparrow$) 1.000 0.065 0.092 0.449 0.769 0.611 1.000 1.000 1.000 0.765
+Identical LLaMA-7B BERTScore ($\uparrow$) 1.000 0.065 0.092 0.449 0.769 0.611 1.000 1.000 1.000 0.765
 ROUGE-L ($\uparrow$) 1.000 0.036 0.062 0.500 0.562 0.436 1.000 1.000 1.000 0.687
@@
 ROUGE-L ($\uparrow$) 0.000 0.000 0.001 0.825 0.443 0.564 0.994 0.994 0.994 0.447
-LLaMA-3.1-8B-Distilled BERTScore ($\uparrow$) 0.083 0.062 0.081 0.642 0.492 0.635 0.894 0.258 0.762 0.610
-ROUGE-L ($\uparrow$) 0.000 0.000 0.001 0.633 0.227 0.413 0.868 0.122 0.479 0.421
 Qwen2.5-Math-7B BERTScore ($\uparrow$) 0.229 0.105 0.105 0.918 0.552 0.783 1.000 0.983 0.996 0.422
 ROUGE-L ($\uparrow$) 0.186 0.000 0.000 0.842 0.355 0.580 1.000 0.977 0.996 0.286
-With this optimization, the online computational overhead is drastically reduced.
-The matrix multiplication with $M$ is now part of the offline weight preparation.
-The online cost for obfuscating and de-obfuscating one KV-cache block is now dominated by the multiplications with $S,\hat{P},S^{-1}$, totaling approximately $b^{3}+2b^{2}d$ floating-point operations.
-For the LLaMA-3.1-8B example ($b=16,d=128,D=4096$), this overhead is merely $0.83\%$ of the cost of re-computing the KV-cache block, rendering the performance impact minimal.
-Detailed analysis can be found in Appendix -F.
-V Evaluation
-V-A Experimental Setup
-Models.
-We first select several state-of-the-art models from the popular LLaMA [9] and Qwen [37] series, including LLaMA-3.2-1B, LLaMA-3.2-3B-Instruct, LLaMA-3.1-8B, Qwen2.5-Math-7B, and DeepSeek-R1-Distill-LLaMA-8B (LLaMA-3.1-8B-Distilled) [10], which is fine-tuned from LLaMA-3.1-8B.
-In addition to these five models that use Grouped-Query Attention (GQA), we also leverage a classic models with Multi-Head Attention (MHA), namely LLaMA-7B [32], to demonstrate the generalizability of our method.
-Datasets.
-To evaluate the effectiveness of our attacks, we use the lmsys-chat-1m dataset [43], which is collected from a real-world inference service and contains a rich variety of dialogues and instructions.
-We randomly sample 1,000 instances from this dataset as the test set for validating attack effectiveness.
-To assess the impact of KV-Cloak on model accuracy and inference latency, we employ standard language model benchmarks, including MMLU [13, 12] and SQuAD [25].
-Attack and Defense Evaluation Metrics.
-BERTScore [41] and ROUGE-L [19] are used to measure the similarity between the original input and the text reconstructed from the KV-cache.
-BERTScore, which is based on the all-mpnet-base-v2 model, is better at capturing semantic similarity, while ROUGE-L primarily reflects lexical-level precision and recall.
-V-B Attack Effectiveness
-In this section, we first evaluate the privacy leakage risk of using the KV-cache by implementing our proposed inversion, collision, and injection attacks.
-Additional experiments, such as the ablation study of attacks, can be found in Appendix -C.
-Experimental Settings.
-We attack 7 LLMs as described in Section V-A.
-Notably, while attacking the LLaMA-3.1-8B-Distilled, which is a fine-tuned version, we assume that the attacker can only get access to its base model (i.e., LLaMA-3.1-8B).
-It can be utilized to validate the attack effectiveness against fine-tuning models.
-Furthermore, the inversion and collision attacks require KV-cache data from only a single layer.
-Given that models have varying numbers of layers, we select the KV-cache from the first, middle, and last layers to perform the inversion and collision attacks.
-In contrast, the injection attack is conducted on the complete KV-cache data, encompassing all layers.
-Results of KV-cache Inversion Attack.
-As shown in Table I, our experimental results generally align with our analysis: a successful inversion attack requires two sufficient conditions to be met: (1) access to the KV-cache from the model’s first layer, and (2) an invertible attention projection matrix, which is characteristic of the MHA mechanism.
-The inversion attack can reconstruct the user input with near-perfect fidelity (approaching 100%) from the first-layer KV-cache of MHA-based models.
-In contrast, reconstruction attempts using the KV-cache from subsequent layers yield completely unintelligible results (mostly <0.1<0.1).
-However, we also find that for the LLaMA-3.2-1B model using GQA, the inversion attack achieved a reconstruction fidelity of nearly 100% on the first-layer KV-cache.
-This may be because the ratio of the rank of this model’s attention projection matrix to its hidden state dimension is particularly high.
-Consequently, even though the projection matrix is not a square, invertible matrix, the least squares method can still recover the input to the attention layer with high accuracy.
-Takeaway 1: Inversion attack is mostly effective against the KV-cache of the first MHA layer.
-Results of KV-cache Collision Attack.
-For the collision attack, we measure the distance between a locally generated KV-cache and the leaked KV-cache using the Frobenius norm.
-We experimentally observed that the distances ($dis_{other}$) between the leaked KV-cache and the caches generated from incorrect tokens in the vocabulary approximate a Gaussian distribution, as shown in Figure 8.
-Consequently, we leverage the statistical properties of this distribution (i.e., its mean and standard deviation) to identify the correct token, whose corresponding distance ($dis_{target}$) will manifest as a statistical outlier.
-After experimental analysis, we defined outliers as values falling below 3 standard deviations from the mean of the $dis_{batch}$ distribution.
-The batch size for local analysis was set to 256.
-Details of the ablation study can be found in Appendix -C1.
-Under these settings, the experimental results for the collision attack are presented in the “Collision” column of Table I, demonstrating high reconstruction accuracy across all layers of all models.
-We also experimented with truncating the probability-sorted vocabulary on the LLaMA-3.2-1B model, aiming to reduce the time lost on unlikely tokens.
-The results are presented in Figure 6.
-We found that even when searching only the top $1/8$ of tokens with the highest predicted probabilities, the average input reconstruction fidelity remains at approximately 96.1% of what is achieved by searching the entire vocabulary.
-Meanwhile, the attack time is reduced to less than 42.9% of the full search, with the average time to reconstruct a user input from a single layer’s KV-cache dropping from 5.06s to 2.17s.
-This confirms that truncating the vocabulary based on model-predicted probabilities is a highly effective strategy for accelerating the collision attack.
-Figure 6: The effect of truncating the probability-sorted vocabulary on reconstruction fidelity and attack time.
-Experiments were run with a batch size of 256 and an outlier threshold of $3\sigma_{other}$.
-The experimental results for the collision attack demonstrate that this method overcomes the limitations of the inversion attack, which is restricted to specific layers (i.e., the first layer) and attention mechanisms (i.e., MHA).
-For open-source models, the collision attack can reconstruct user inputs with high accuracy and acceptable efficiency from the KV-cache of any layer.
-Crucially, it is also effective against the LLaMA-3.1-8B-Distilled model—a model fine-tuned from Meta-LLaMA-3.1-8B—by using the public weights of its open-source base model.
-Takeaway 2: Collision attack achieves relatively high accuracy, demonstrating its universality and the high-risk threat it poses to real-world systems.
-Collision Attack Enhanced with Prior Knowledge (Collision+).
-As shown in Table I, our enhanced collision attack (assuming rank $r=8$) is highly potent, achieving near-perfect reconstruction accuracy (approaching 100%) on any layer of any tested open-source model, while also demonstrating notable effectiveness against fine-tuned models.
-We also conducted experiments on the LLaMA-3.1-8B-Distilled model using optimal thresholds derived for different assumed ranks ($r$).
-The results are shown in Figure 7.
-For the first layer, the actual average rank of the target token is approximately 4; using the optimal threshold for this rank increases the reconstruction accuracy to 138.6% of the unenhanced attack.
-For the middle layer, the actual rank appears to be greater than 128, as none of the tested enhanced thresholds surpassed the accuracy of the baseline attack.
-For the last layer, the optimal rank is between 64 and 128, and applying the corresponding enhanced threshold boosts the reconstruction accuracy to approximately 139.0% of the baseline.
-Figure 7: Collision attack experiments on LLaMA-3.1-8B-Distilled using optimal thresholds derived for different assumed token ranks ($r$).
-Takeaway 3: Augmenting the collision attack with model-specific prior knowledge achieves near-perfect ($\approx 100\%$) user-input recovery accuracy.
+Finetune LLaMA-3.1-8B-Distilled BERTScore ($\uparrow$) 0.083 0.062 0.081 0.642 0.492 0.635 0.894 0.258 0.762 0.610
+(LLaMA-3.1-8B) ROUGE-L ($\uparrow$) 0.000 0.000 0.001 0.633 0.227 0.413 0.868 0.122 0.479 0.421
+Cross-Arch LLaMA-2-7B BERTScore ($\uparrow$) 0.067 0.086 0.084 0.058 0.070 0.069 0.075 0.062 0.060 0.087
+(LLaMA-7B) ROUGE-L ($\uparrow$) 0.038 0.057 0.031 0.013 0.000 0.002 0.052 0.017 0.024 0.019
+This section evaluates the feasibility of reconstructing user inputs via the proposed Inversion, Collision, and Injection attacks.
+We systematically test these vectors against the models defined in Section IV-A.
+Detailed ablation studies are provided in Appendix -C.
+IV-B1 Experimental Settings
+Beyond standard reconstruction, we introduce two specific scenarios to evaluate attack robustness under restricted adversarial knowledge: (1) Fine-tuning Mismatch: We attack the fine-tuned LLaMA-3.1-8B-Distilled assuming the adversary only possesses weights from its base model (LLaMA-3.1-8B).
+This validates effectiveness when exact model weights are unavailable.
+(2) Cross-Architecture Mismatch: We employ LLaMA-7B parameters to attack LLaMA-2-7B to assess generalizability across disjoint architectures.
+Regarding data scope, the Inversion and Collision attacks operate on single-layer KV-caches.
+To characterize layer-wise vulnerability, we evaluate these attacks on the first, middle, and last layers respectively.
+In contrast, the Injection Attack utilizes the complete, multi-layer KV-cache to leverage the model’s full semantic processing capabilities.
+IV-B2 Results of KV-cache Inversion Attack
+As detailed in Table I, the efficacy of the Inversion Attack generally hinges on two conditions: (1) access to the first-layer KV-cache, and (2) an algebraically invertible projection matrix, typical of MHA.
+Under these constraints, we achieve near-perfect reconstruction fidelity ($\approx 100\%$), whereas deeper layers yield unintelligible noise ($10\%$) due to semantic fusion.
+Notably, the GQA-based LLaMA-3.2-1B also exhibits high first-layer vulnerability ($\approx 100\%$).
+We attribute this to the model’s high projection rank relative to its hidden state dimension.
+This distinct property allows the least squares method to effectively recover inputs despite the non-square nature of the projection matrix.
+Takeaway 1: The Inversion Attack is effective against the first-layer KV-cache of MHA models, with exceptions for architectures possessing high-rank projection matrices.
+IV-B3 Results of KV-cache Collision Attack
 (a) The distributions of LLaMA-3.1-8B model.
 (b) The distributions of LLaMA-3.1-8B-Distilled model.
-Figure 8: Distributions of $dis_{target}$ (orange) and $dis_{other}$ (blue) from a collision attack.
-The input is an excerpt from “The Bitter Lesson”.
-The attack uses the LLaMA-3.1-8B model’s weights against the last-layer KV-cache generated by itself (left) and by LLaMA-3.1-8B-Distilled (right).
-The green dotted line shows the baseline heuristic threshold ($3\sigma_{other}$), while the red dashed line shows the enhanced threshold derived from prior knowledge (assuming rank $r=64$).
-To further illustrate this, we analyzed the distance distributions using an excerpt from “The Bitter Lesson” as input.
-Figure 8 shows these distributions for an attack on the last-layer KV-cache.
-We observe that for the open-source model (left panel), the baseline heuristic threshold of $3\sigma_{other}$ results in a 0.13% false-positive rate (i.e., 0.13% of $dis_{other}$ distances are misclassified).
-If we assume the correct token is at rank $r=64$ in the sorted vocabulary, this yields a per-token attack success rate of 91.84%.
-In contrast, by leveraging a more precise threshold derived from statistical analysis, this success rate can be boosted to a perfect 100%.
-For the fine-tuned model (right panel), the enhanced threshold improves the per-token success rate from 91.79% to 97.82%.
-This substantial improvement in per-token accuracy leads to a dramatic increase in the overall reconstruction fidelity for the entire input sequence.
-Results of KV-cache Injection Attack.
-The success of this attack hinges on two factors: the model’s ability to effectively parse the semantic content of the KV-cache and its comprehension of the injected instruction.
-For our evaluation, we employed the most effective instruction identified in our ablation study, “Repeat the previous content.” (the selection process is detailed in Appendix -C2), to attack the KV-cache data from each target model.
-As shown in Figure I, the injection attack yielded an average BERTScore of 0.58 and a ROUGE-L score of 0.42 for the reconstructed user inputs.
-These results are consistently lower than those achieved by the more direct collision attack.
-The attack performed best on the LLaMA-7B.
-This supports the hypothesis that the MHA mechanism preserves more contextual information than the more compressed attention variants (e.g., GQA) within the KV-cache.
-Takeaway 4: Absent verbatim recovery, injection attacks can still exfiltrate the core information of user inputs through semantic-level interaction.
-The reconstruction accuracy of the injection attack is generally lower than that of the collision attack, and it requires access to the complete KV-cache from all layers.
-However, its key advantage is speed: the attack’s execution time is equivalent to a single inference step, making it significantly faster than the iterative collision attack.
-Moreover, the existence of this attack vector imposes an additional requirement on any defense mechanism: it must ensure that a protected KV-cache is rendered unintelligible even to a standard, unprotected LLM.
-V-C Evaluation of KV-Cloak
-V-C1 Experimental Settings
-In order to comprehensively evaluate our proposed KV-Cloak scheme, this section presents a systematic comparison against two baselines: (1) a standard, unprotected system (Plaintext) and (2) a defense based on differential privacy with Gaussian noise (DP).
-Our evaluation is structured around three core dimensions: Security, Model Accuracy, and Performance Overhead.
-The results demonstrate that KV-Cloak is the only solution that provides robust security protection with acceptable performance overhead, all without sacrificing model accuracy.
-In our KV-Cloak, the obfuscation involves matrix multiplication of KV-cache blocks with invertible secret matrices $S$ and $M$, along with the addition of a random mask $A$.
-  • Block Size: To ensure compatibility with the PagedAttention mechanism used in modern inference engines like vLLM, we experimented with common block sizes: 16, 32, and 64.
-  • Invertible Matrices $S,M$: To avoid precision loss from matrix inversion during de-obfuscation, we sample $S$ and $M$ from the space of orthogonal matrices.
-  • Additive Mask $A$ and Padding: The values in the mask $A$ and the padding for incomplete blocks are kept small to minimize their impact on numerical precision.
-We sample the elements of $A$ from the range $3\theta_{K},4\theta_{K}]$ and use a value of $1.5\theta_{K}$ for padding.
-Here, $\theta_{K}$ is the maximum absolute value observed in the K-cache elements when inferring on an excerpt from “The Bitter Lesson”.
-The same methodology is applied for the V-cache.
-To establish a robust DP baseline, we conducted an ablation study (detailed in Appendix -D) to select a configuration that balances utility and privacy for comparison against KV-Cloak.
-Based on our findings, we selected a practical configuration: the clipping threshold was set to the 50th percentile of the L2 norm distribution observed across the dataset.
-Subsequently, we applied Gaussian noise calibrated to satisfy ($\epsilon=10^{8},\delta=10^{-5}$)-DP independently to both the K-cache and V-cache.
-V-C2 Evaluation of Security
-TABLE II: Efficacy of Defense Mechanisms Against Input Reconstruction Attacks.
+Figure 3: Distance distributions of target tokens $d_{\text{target}}$ (orange) versus incorrect tokens $d_{\text{other}}$ (blue) in the Collision Attack.
+The input is an excerpt from “The Bitter Lesson”(see Appendix -B).
+The attack targets the last-layer KV-cache of LLaMA-3.1-8B (left) and LLaMA-3.1-8B-Distilled (right) using weights from the base model.
+Vertical lines indicate the heuristic threshold ($3\sigma_{\text{other}}$, green dotted) and the prior-knowledge enhanced threshold ($r=64$, red dashed).
+Figure 4: The effect of truncating the probability-sorted vocabulary on reconstruction fidelity and attack time.
+Experiments were run with a batch size of 256 and an outlier threshold of $3\sigma_{\text{other}}$.
+Unlike the Inversion Attack, the Collision Attack relies on forward-pass matching using the Frobenius norm.
+As shown in Figure 3, we observe that distances for incorrect tokens ($d_{\text{other}}$) approximate a Gaussian distribution, whereas the target token ($d_{\text{target}}$) manifests as a distinct statistical outlier.
+By setting an outlier threshold of $3\sigma$ below the batch mean (batch size=256, see Appendix -C1), we achieve high reconstruction accuracy across all layers of all tested models (Table I).
+Crucially, this method overcomes the architectural constraints of MHA; it is effective even against the fine-tuned LLaMA-3.1-8B-Distilled using only public base model weights, demonstrating robustness against parameter divergence.
+Efficiency via Probability-Guided Pruning.
+To enhance practicality, we evaluate truncating the search space based on model-predicted probabilities.
+Experimental results on LLaMA-3.2-1B (Figure 4) indicate that searching only the top $1/8$ of tokens retains 96.1% of the full-search fidelity.
+This optimization reduces the average reconstruction time per layer from 5.06s to 2.17s ($43\%$ of the original latency), rendering the attack highly efficient for real-time scenarios.
+Takeaway 2: The Collision Attack is a universal threat, achieving high fidelity across diverse architectures (including fine-tuned models) and layers, with high efficiency via probability pruning.
+IV-B4 Collision Attack Enhanced with Prior Knowledge (Collision+)
+We further evaluate the potency of integrating adversarial prior knowledge (i.e., adaptive thresholding based on assumed token rank $r$).
+Figure 5: Collision Attack experiments on LLaMA-3.1-8B-Distilled using optimal thresholds derived for different assumed token ranks ($r$).
+Efficacy on Fine-tuned Models.
+As detailed in Table I, assuming a rank $r=8$ yields near-perfect reconstruction ($\approx 100\%$) across all open-source models.
+For the fine-tuned LLaMA-3.1-8B-Distilled, we analyze the impact of rank assumptions in Figure 5.
+In the first layer, where the actual average target rank is $\approx 4$, using the optimal threshold boosts accuracy to 138.6% of the baseline.
+Conversely, the middle layer shows no improvement, implying an actual rank >128>128.
+In the last layer, optimal rank alignment ($64\leq r\leq 128$) increases accuracy to 139.0% relative to the baseline.
+Statistical Justification.
+To further illustrate this, we analyze distance distributions using “The Bitter Lesson” excerpt (Figure 3).
+For the open-source model, the heuristic $3\sigma_{\text{other}}$ threshold yields a 0.13% false-positive rate, capping per-token success at 91.84% (assuming $r=64$).
+By switching to a statistically derived threshold, we eliminate false positives, achieving 100% success.
+Similarly, for the fine-tuned model, the enhanced threshold improves per-token success from 91.79% to 97.82%, dramatically increasing full-sequence reconstruction fidelity.
+Takeaway 3: Augmenting the Collision Attack with model-specific prior knowledge achieves near-perfect ($\approx 100\%$) user-input recovery accuracy.
+IV-B5 Results of KV-cache Injection Attack
+We evaluate semantic exfiltration using the optimal directive identified in Appendix -C2: “Repeat the previous content”.
+As detailed in Table I, this vector achieves an average BERTScore of 0.58 and ROUGE-L of 0.42.
+While these metrics are lower than those of the exact-match Collision Attack, they confirm significant leakage of the input’s core meaning.
+Notably, the attack exhibits peak performance on LLaMA-7B.
+This supports the hypothesis that the MHA mechanism retains higher contextual fidelity within the KV-cache compared to compressed variants like GQA.
+Consequently, MHA caches are more semantically intelligible to the model, increasing vulnerability to instruction-based extraction.
+This imposes a critical requirement for defenses: a protected KV-cache must be rendered semantically unparsable to the LLM itself, preventing the model from being weaponized to interpret stolen contexts.
+Takeaway 4: Even without verbatim recovery, Injection Attacks successfully exfiltrate core user intent, necessitating defenses that neutralize semantic intelligibility.
+IV-C Attack Robustness under Partial Knowledge Scenarios
+To assess practical applicability, we evaluate the robustness of our attacks under a relaxed threat model characterized by two realistic constraints: incomplete data interception and model parameter mismatch.
+Incomplete KV-cache Data.
+Our analysis reveals distinct data requirements across attack vectors.
+The Injection Attack necessitates the full, cross-layer context and fails when data is fragmented.
+Conversely, the Collision Attack exhibits superior robustness; possessing the KV-cache from any single layer is sufficient to perform high-fidelity input reconstruction.
+Model Mismatch.
+We further evaluate the scenario where the adversary’s local model differs from the target service model.
+Experiments confirm that attack effectiveness hinges on the correlation of weight parameters.
+  • Fine-tuning Mismatch (Gray-box): We targeted a fine-tuned model (LLaMA-3.1-8B-Distilled) assuming the adversary only possesses the open-source base model (LLaMA-3.1-8B).
+Results in Table I show that the attack remains highly effective, suggesting that fine-tuning preserves the fundamental weight correlations exploited by our mechanism.
+  • Architectural Mismatch (Black-box): In contrast, cross-architecture attacks (e.g., using LLaMA-7B to attack LLaMA-2-7B) fail completely.
+Reconstruction similarity drops to levels statistically indistinguishable from random guessing (<0.1<0.1).
+This confirms that the Collision Attack relies on the algebraic correspondence of weight parameters, which is absent across disjoint architectures.
+V KV-Cloak: A Lightweight KV-cache Defense
+V-A Motivation for KV-Cloak
+Limitations of Existing Privacy-Preserving Techniques.
+Current privacy-preserving techniques fundamentally struggle to balance security, efficiency, and utility when applied to the massive, latency-sensitive KV-cache.
+We analyze their specific limitations below:
+  • Cryptographic Methods (Overhead Bottleneck): Standard cryptographic techniques [4, 34], such as symmetric AES encryption, provide confidentiality by ensuring data remains in ciphertext during storage and transmission, requiring decryption only for legitimate use.
+A more advanced approach, Homomorphic Encryption (HE) [26, 2], allows for direct computation on encrypted data, theoretically enabling parts of the attention mechanism to operate on the KV-cache without ever decrypting it.
+However, despite their strong security guarantees, the immense computational overhead and latency introduced by these methods are prohibitive.
+Given that the KV-cache can be tens or hundreds of gigabytes, applying full encryption or HE is unsuitable for the high-throughput, low-latency requirements of LLM inference.
+  • Differential Privacy (Utility Trade-off): Applying Differential Privacy involves injecting calibrated noise into the Key-Value vectors to mask individual data points [1, 20, 42].
+However, the sparse and sensitive nature of attention mechanisms creates a severe utility-privacy trade-off.
+To achieve a meaningful level of privacy, the required amount of noise would significantly degrade the LLM’s inference accuracy to an unacceptable degree.
+  • KV-Shield (Security and Compatibility Flaws): KV-Shield [41] is the only existing lightweight obfuscation scheme for KV-cache.
+It synchronously permutes attention weight rows ($W_{q},W_{k},W_{v}$) to shuffle cache elements, obfuscating them before an attention score is calculated.
+The obfuscated attention output is then de-obfuscated for subsequent steps.
+We identify two critical failures in this design: (1) Security Flaws.
+First, the simple shuffling preserves the inherent statistical distribution of the data, leaving it highly vulnerable to our proposed Collision Attacks.
+Second, the reliance on a fixed obfuscation key lacks dynamic randomness, rendering the scheme defenseless against CPA.
+(2) Architectural Incompatibility.
+The shuffling disrupts relative positioning, making it incompatible with RoPE used in SOTA models (e.g., LLaMA, Qwen).
+Design Objectives. The deficiencies of prior works necessitate a specialized defense for KV-cache. We define three mandatory design goals for a practical solution:
+  • Robust Security: The mechanism must resist targeted reconstruction attacks by effectively masking both the algebraic and statistical properties of the cache.
+  • Lossless Model Fidelity: It must preserve the exact mathematical equivalence of the attention mechanism, ensuring zero degradation in generation quality.
+  • Negligible Overhead: The defense must operate with minimal latency, ideally shifting computational costs offline to maintain high inference throughput.
+To meet these objectives, we propose KV-Cloak, a novel defense mechanism that synergizes reversible matrix obfuscation with operator fusion to deliver robust security, lossless model fidelity, and negligible inference overhead.
+V-B Naive Defense: Reversible Linear Obfuscation
+To mitigate privacy risks with minimal overhead, we first consider a naive defense that employs reversible linear transformations to obscure the statistical properties of the KV-cache.
+In the context of modern inference frameworks (e.g., vLLM), we denote the key vectors for a single attention head within a PagedAttention block as a matrix $K\in\mathbb{R}^{b\times d}$, where $b$ is the block size and $d$ is the head dimension.
+Since the protection mechanism applies analogously to value vectors, we focus our analysis on $K$.
+The obfuscation transformation is defined as:
+$K^{\prime}=SKM,$ (7)
+where $S\in\mathbb{R}^{b\times b}$ and $M\in\mathbb{R}^{d\times d}$ are secret, randomly generated invertible matrices.
+This transformation aims to conceal the raw content of $K$ while preserving the matrix dimensions necessary for storage allocation.
+Security Analysis.
+Despite its operational simplicity, this scheme is fundamentally insecure under a CPA model due to its fixed linear structure.
+While a real-world adversary cannot generate arbitrary matrix $K$ directly (as $K$ is constrained by the model’s embedding projection of valid tokens), we demonstrate that this constraint does not preclude a full compromise.
+An adversary can circumvent this restriction by mounting a differential attack through carefully crafted prompts.
+Specifically, the adversary chooses two inputs yielding plaintexts $K_{1}$ and $K_{2}$, thereby controlling the difference $\Delta K=K_{1}-K_{2}$.
+Due to the linearity of the scheme, the observed ciphertext difference is $\Delta K^{\prime}=S(\Delta K)M$.
+By systematically crafting inputs such that $\Delta K$ approaches a series of standard basis matrices (i.e., matrices with a single non-zero entry), the adversary can isolate and solve for the columns of $S$ and the rows of $M$.
+This algebraic attack enables full recovery of the secret matrices (up to a scalar ambiguity) with a computational complexity of only $O(b^{2}d+bd^{2})$, rendering the naive defense ineffective against determined adversaries.
+V-C One-Time Pad Block-wise Shuffling
+Figure 6: Illustration of the KV-Cloak block-wise shuffling mechanism.
+By permuting $K$ and $V$ vectors within a block, the mechanism eliminates positional side channels while maintaining mathematical equivalence in the attention output.
+V-C1 Eliminating Redundant Positional Information in KV-cache
+To enhance the security of the obfuscation scheme, we introduce additional randomness by eliminating the architectural redundancy of the KV-cache.
+We observe that the physical memory ordering of $k,v$ vectors is computationally superfluous during inference, as positional semantics are intrinsic to the vectors via RoPE.
+Mechanism.
+We implement a dynamic, block-wise random permutation.
+By reordering $k,v$ pairs within each block while maintaining their internal correspondence, we decorrelate the physical storage index from the logical token sequence.
+As illustrated in Figure 6, this operation preserves the mathematical invariance of the attention mechanism while imposing a combinatorial complexity barrier of $b!$ on the adversary.
+Given typical block sizes $b\in\{16,32,64\}$ [17], this renders brute-force matching computationally infeasible.
+Formalization. The enhanced obfuscation employs a locally generated, one-time pad permutation matrix $\hat{P}$ before applying the linear transformations:
+$K^{\prime}=S\hat{P}KM.$ (8)
+Crucially, for efficiency, $\hat{P}$ is ephemeral and does not require storage.
+The inference engine performs subsequent computations directly on the de-obfuscated permuted state $\hat{P}K$ without reconstructing the original order, thereby incurring zero additional storage overhead for the permutation key.
+V-C2 Rank Preservation and Implicit Key Recovery
+Figure 7: Value distribution of KV-cache elements ($K$ and $V$) for LLaMA-3.2-1B. The data is collected during inference on “The Bitter Lesson”.
+A critical challenge arises when the original matrix $K$ exhibits low rank (e.g., rank $1$ in cases of identical repeated tokens).
+In such scenarios, the permutation entropy collapses, as row shuffling becomes statistically invisible.
+To guarantee obfuscation robustness, we introduce a secret additive mask matrix $A$ before permutation.
+Thereby, we can ensure the transformed matrix $(K+A)$ maintains sufficient rank to preserve cryptographic strength of $\hat{P}$.
+However, explicitly storing the unique one-time pad $\hat{P}$ for every block would reintroduce significant storage overhead.
+We resolve this by designing $A$ to enable implicit key recovery without $\hat{P}$.
+Magnitude-based Positional Embedding.
+We exploit the numerical characteristics of attention activations.
+As shown in Figure 7, elements of $k,v$ vectors are statistically bounded by a threshold $\theta_{K}$ (e.g., typically <100<100).
+Leveraging this, we construct $A$ to function as a “positional beacon” by embedding values significantly larger than $\theta_{K}$ into specific coordinates.
+This ensures that the matrix $(K+A)$ maintains enough rank.
+Zero-Storage Recovery.
+Since the “beacons” in $A$ dominate the magnitude of $K$, the permuted mask term $\hat{P}A$ remains statistically separable from $\hat{P}K$ within the ciphertext $\hat{P}(K+A)$.
+During de-obfuscation, the system dynamically identifies the row permutation indices by detecting these high-magnitude outliers, allowing it to reconstruct $\hat{P}A$ on-the-fly.
+Subtracting this reconstructed mask yields the permuted state $\hat{P}K$ required for inference.
+This mechanism effectively offloads the storage of $\hat{P}$ into the data itself.
+Formalization. The complete obfuscation logic integrates this additive masking with the linear and permutation layers:
+$K^{\prime}=S\hat{P}(K+A)M,$ (9)
+[…]
+V-C3 Security Analysis
+The security of KV-Cloak relies fundamentally on the confidentiality of its secret matrices ($S,M,A$).
+Under a standard CPA model, an adversary attempting to recover these keys faces a prohibitive computational barrier.
+Our analysis indicates that the time complexity to solve for the keys is $O((b^{2}d+bd^{2})\cdot b!)$.
+Crucially, the factorial term $b!$, introduced by the dynamic block-wise permutation, renders brute-force key recovery computationally infeasible.
+This computational hardness is complemented by layered security mechanisms designed to mitigate specific attack vectors: (1) The one-time pad permutation severs the explicit token-to-vector positional correspondence at an information-theoretic level, invalidating the premise of Collision Attacks.
+(2) A reversible algebraic transformation via secret matrices ($S,M$) completely disrupts the cache’s statistical distribution properties (e.g., mean, variance), neutralizing profiling and data mining attacks.
+(3) The combined obfuscation renders the cache semantically unintelligible and unparsable by pristine (unmodified) models, thereby thwarting Injection Attacks that attempt to leverage the model’s own capabilities.
+V-D Implicit Obfuscation via Operator Fusion
+To minimize the runtime latency overhead introduced by the obfuscation and de-obfuscation processes, we propose an operator fusion strategy.
+This approach integrates the obfuscation operators directly into the model’s weight matrices offline, thereby rendering the online computational impact negligible while maintaining strict mathematical equivalence.
+Transformed Vectors and Invariance.
+We introduce two secret, invertible matrices, $M_{1},M_{2}\in\mathbb{R}^{d\times d}$, to transform the query, key, and value vectors ($q,k,v$) and the output projection $W_{o}$.
+The transformed vectors ($q^{m},k^{m},v^{m}$) and the modified output weight $W_{o}^{m}$ are defined as:
+$\left\{\begin{array}[]{l}q^{m}=q(M_{1}^{-1})^{\top},\\ k^{m}=kM_{1},\\ v^{m}=vM_{2},\\ W_{o}^{m}=W_{o}(M_{2}^{-1})^{\top}.\end{array}\right$. (10)
+This design ensures that the core attention mechanism remains invariant. Specifically, the attention scores (scaled dot-product of query and key) are preserved:
+$q_{i}^{m}(k_{j}^{m})^{\top}$ $=\left(q_{i}(M_{1}^{-1})^{\top}\right)\left(k_{j}M_{1}\right)^{\top}=q_{i}(M_{1}^{-1})^{\top}M_{1}^{\top}k_{j}^{\top}$ (11)
+$=q_{i}(M_{1}M_{1}^{-1})^{\top}k_{j}^{\top}=q_{i}k_{j}^{\top}.$
+Similarly, the output of the attention head is mathematically identical to the unprotected, guaranteeing lossless accuracy:
+$v_{j}^{m}(W_{o}^{m})^{\top}$ $=(v_{j}M_{2})\left(W_{o}(M_{2}^{-1})^{\top}\right)^{\top}$ (12)
+$=v_{j}M_{2}(M_{2}^{-1})W_{o}^{\top}=v_{j}W_{o}^{\top}.$
+This invariance guarantees that the model’s output is mathematically identical to that of the original, unprotected model, ensuring lossless accuracy.
+Fusion under RoPE. Implementing this fusion requires incorporating the input projection $x$ and the position-dependent RoPE matrix $R_{\Theta,i}^{d}$. The transformed vectors are expressed as:
+$\left\{\begin{array}[]{l}q^{m}=xW_{q}^{\top}R_{\Theta,i}^{d}(M_{1}^{-1})^{\top}\\ k^{m}=xW_{k}^{\top}R_{\Theta,i}^{d}M_{1}\\ v^{m}=xW_{v}^{\top}M_{2}\\ W_{o}^{m}=W_{o}(M_{2}^{-1})^{\top}\end{array}\right.$.
+(13)
+From Eq.
+13, fusing $M_{2}$ into the value and output weights is straightforward: we pre-compute $(W_{v}^{m})^{\top}=W_{v}^{\top}M_{2}$ and $W_{o}^{m}=W_{o}(M_{2}^{-1})^{\top}$.
+However, fusing $M_{1}$ into $W_{k}$ and $W_{q}$ presents a challenge due to the intermediate application of RoPE.
+The fusion is algebraically feasible if and only if the secret matrix $M_{1}$ and the RoPE matrix $R_{\Theta,i}^{d}$ commute, i.e., $R_{\Theta,i}^{d}M_{1}=M_{1}R_{\Theta,i}^{d}$.
+In this case, Eq.
+13 simplifies to:
+$k^{m}=xW_{k}^{\top}(M_{1}R_{\Theta,i}^{d})=x(W_{k}^{\top}M_{1})R_{\Theta,i}^{d}.$ (14)
+This reordering allows us to absorb $M_{1}$ into the weight matrix as $(W_{k}^{m})^{\top}=W_{k}^{\top}M_{1}$. A symmetric logic applies to $W_{q}$.
+To satisfy this commutativity constraint, we structurally design $M_{1}$ as a invertible block-diagonal matrix composed of $2\times 2$ rotation-scaling sub-matrices, analogous to the structure of RoPE (detailed in Appendix -A).
+Additionally, we constrain $M_{2}$ to be a random invertible rotation-scaling matrix and calibrate the scaling factors of both matrices.
+This ensures that after the transformation with mask $A$, the padding tokens remain statistically distinguishable as outliers for implicit key recovery.
+Offline Computation and Online Efficiency. Consequently, the obfuscation operators are fully fused into the attention layer parameters offline. The new pre-computed weights are:
+$\left\{\begin{aligned} W_{q}^{m}&=M_{1}^{-1}W_{q}\\ W_{k}^{m}&=M_{1}^{\top}W_{k}\\ W_{v}^{m}&=M_{2}^{\top}W_{v}\\ W_{o}^{m}&=W_{o}(M_{2}^{-1})^{\top}\end{aligned}\right.$. (15)
+During online inference, the KV-Cloak obfuscation is applied to the cache generated by these fused weights:
+$K^{\prime}=S\hat{P}(K^{m}+A),$ (16)
+By eliminating explicit online matrix multiplications with $M$, the cost for protecting one KV-cache block is reduced to the multiplications with $S$, $\hat{P}$, and $S^{-1}$.
+This totals approximately $b^{3}+2b^{2}d$ floating-point operations.
+For an LLaMA-3.1-8B instance ($b=16,d=128,D=4096$), this overhead represents merely 0.83% of the KV-cache re-computation cost (detailed in Appendix -F), verifying that KV-Cloak achieves robust security with minimal performance impact.
+TABLE II: Comparison of defense mechanisms (KV-Cloak vs. DP) against input reconstruction attacks on LLaMA-7B, LLaMA-3.2-1B, and LLaMA-3.1-8B-Distilled.
 Model Protect Type Metric Inversion Collision Collision+ Injection
@@
 ROUGE-L ($\downarrow$) 0.003 0.291 0.185 0.351 0.530 0.122 0.543 0.049
-This section evaluates the capability of different protection schemes to defend against our three KV-cache-based reconstruction attacks.
-We repeat the attack experiments on the KV-cache after applying each defense mechanism.
-As shown in Table II (evaluation on the remaining models can be found in Appendix -E), for the “Plaintext” baseline, all three attacks achieve very high success rates.
-After applying KV-Cloak, the semantic similarity (BERTScore) of all attack outputs drops to a level consistent with random chance, and the ROUGE-L score falls to nearly 0.
-These results are statistically indistinguishable from comparing the original input with a random string, demonstrating that semantic reconstruction is entirely disrupted.
-This proves that KV-Cloak effectively protects the private information within the KV-cache.
-For the DP baseline, its defensive efficacy is strongly correlated with the privacy budget $\epsilon$.
-With a weak budget of $\epsilon=10^{8}$, the accuracy of the inversion and injection attacks is reduced, but the collision attack can still recover some useful information.
-We also compare the distributions of $dist_{target}$ and $dist_{other}$ for the collision attack on the last layer of LLaMA-3.2-1B’s KV-cache under different protections.
-As shown in Figure 9, the two distributions remain effectively distinguishable under DP protection.
-When using an enhanced threshold derived from prior knowledge, the per-token attack success rate is as high as 94.84% for ($10^{7},10^{-5}$)-DP, and achieving nearly 100% (the same as plaintext) for ($10^{7},10^{-5}$)-DP.
-In contrast, under KV-Cloak’s protection, the two distributions become completely indistinguishable, resulting in a 0% attack success rate.
+VI Evaluation of KV-Cloak
+VI-A Experimental Settings
+We perform a systematic evaluation across three dimensions: Security, Model Accuracy, and Performance Overhead.
+To benchmark the efficacy of KV-Cloak, we compare it against three distinct baselines: (1) a standard, unprotected system (Plaintext), (2) a defense based on differential privacy with Gaussian noise (DP), and (3) a defense via standard cryptographic encryption (AES).
+KV-Cloak Configuration.
+Our implementation involves matrix multiplications with invertible secret matrices $S$ and $M$, and an additive mask $A$.
+We configure these parameters to balance security and numerical stability:
+  • Block Size $b$: To maintain compatibility with the PagedAttention mechanism (e.g., vLLM), we experimented with standard block sizes of $b\in\{16,32,64\}$.
+  • Secret Matrices $S,M$: To minimize precision loss during the matrix inversion required for de-obfuscation, we sample $S$ and $M$ strictly from the orthogonal group.
+  • Additive Mask $A$ and Padding: To preserve numerical precision, the additive mask $A$ and padding values are magnitude-constrained.
+We sample elements of $A$ uniformly from $3\theta_{K},4\theta_{K}]$ and set padding to $1.5\theta_{K}$.
+Here, $\theta_{K}$ denotes the maximum absolute value observed in the K-cache during calibration on an excerpt from “The Bitter Lesson”.
+This methodology is symmetrically applied to the V-cache.
+DP Baseline Configuration.
+To establish a strong DP baseline, we conducted an ablation study (detailed in Appendix -D) to optimize the trade-off between privacy and utility.
+We set the clipping threshold to the 50th percentile of the L2 norm distribution observed across the dataset.
+Based on this, we apply Gaussian noise calibrated for $(\epsilon=10^{8},\delta=10^{-5})$-DP independently to the K/V caches.
+We explicitly select $\epsilon=10^{8}$ because stricter privacy budgets (e.g., $\epsilon=10^{7}$) resulted in near-zero inference accuracy (see Table VIII).
+VI-B Evaluation of Security
 (a) The distributions of LLaMA-3.2-1B model without protection.
@@
 (d) The distributions of LLaMA-3.2-1B model with ($10^{8},10^{-5}$)-DP.
-Figure 9: Distributions of $dis_{target}$ (orange) and $dis_{other}$ (blue) for a collision attack on the last-layer KV-cache.
-The input is an excerpt from “The Bitter Lesson” to the LLaMA-3.2-1B model.
-The distributions are shown under four conditions: (a) Plaintext, (b) KV-Cloak protection, (c) ($10^{7},10^{-5}$)-DP protection, and (d) ($10^{8},10^{-5}$)-DP protection.
-The green dotted line is the baseline heuristic threshold ($3\sigma_{other}$), and the red dashed line is the enhanced threshold assuming rank $r=8$.
-To assess the robustness of our defense, we also tested KV-Cloak’s performance against the enhanced collision attack on multiple LLMs.
-As illustrated in Table II, our mechanism demonstrates strong defensive capabilities.
-With KV-Cloak enabled, the reconstruction accuracy of the attack drops to near-zero, and the recovered outputs are qualitatively equivalent to random noise, confirming that KV-Cloak effectively neutralizes this advanced attack vector.
-Takeaway 5: KV-Cloak completely thwarts all our proposed attacks, reducing the quality of any reconstructed text to a level statistically indistinguishable from random noise.
-V-C3 Inference Accuracy
-To evaluate the model fidelity of our KV-Cloak, we simulate an inference service with a prefill-decode architecture.
-The KV-cache generated during the prefill phase is protected using either DP or KV-Cloak.
-This protected cache is then passed to the decode node (after de-obfuscation in KV-Cloak’s case) for subsequent token generation.
-We then evaluate this protected inference service on the MMLU and SQuAD benchmarks to measure the impact of each method on model accuracy.
+Figure 8: Distance distributions of target tokens $d_{\text{target}}$ (orange) versus incorrect tokens $d_{\text{other}}$ (blue) in the Collision Attack.
+The experiment targets the last-layer KV-cache of LLaMA-3.2-1B using an excerpt from “The Bitter Lesson”.
+Subplots display distributions under four conditions: (a) Plaintext, (b) KV-Cloak protection, (c) ($10^{7},10^{-5}$)-DP protection, and (d) ($10^{8},10^{-5}$)-DP protection.
+Vertical lines indicate the heuristic threshold ($3\sigma_{\text{other}}$, green dotted) and the prior-knowledge enhanced threshold ($r=64$, red dashed).
+This section evaluates the capability of KV-Cloak against the three proposed reconstruction attacks.
+We benchmark against the Plaintext baseline and Differential Privacy mechanisms, applying the attacks to the protected KV-cache.
+Defense Efficacy Comparison.
+As summarized in Table II (evaluation on the remaining models can be found in Appendix -E), the Plaintext baseline is highly vulnerable, yielding high attack success rates on all three attacks.
+In stark contrast, applying KV-Cloak drastically degrades reconstruction quality.
+The BERTScore of all attack outputs drops to a level consistent with random chance, and the ROUGE-L score falls to nearly 0.
+These results are statistically indistinguishable from comparing the original input with a random string, demonstrating that semantic reconstruction is entirely disrupted.
+This proves that KV-Cloak effectively protects the private information within the KV-cache.
+For the DP baseline, security is heavily dependent on the privacy budget $\epsilon$.
+With a weak budget of $\epsilon=10^{8}$, while Inversion and Injection attacks are mitigated, the Collision Attack remains effective, recovering substantial private information.
+Statistical Indistinguishability Analysis.
+To understand the root cause of defense failure or success, we analyze the distance distributions of target tokens ($d_{\text{target}}$) versus other tokens ($d_{\text{other}}$) for the Collision Attack on LLaMA-3.2-1B (Fig.
+8).
+Under DP protection, the two distributions remain statistically distinguishable.
+Consequently, utilizing an enhanced threshold with prior knowledge yields a per-token success rate of 94.84% for ($10^{7},10^{-5}$)-DP, rising to nearly 100% (equivalent to Plaintext) for ($10^{8},10^{-5}$)-DP.
+Conversely, KV-Cloak achieves distribution collapse: the distributions of $d_{\text{target}}$ and $d_{\text{other}}$ become completely indistinguishable.
+This eliminates the statistical separability required for the attack, resulting in a $0\%$ success rate.
+Robustness against Enhanced Attacks.
+To assess the robustness of our defense, We further evaluate by subjecting KV-Cloak to the enhanced Collision Attack (utilizing adversarial prior knowledge) across multiple LLMs.
+As shown in Table II, KV-Cloak demonstrates consistent resilience.
+Even against this advanced vector, reconstruction accuracy remains near-zero, and outputs are qualitatively equivalent to random noise.
+This confirms that KV-Cloak effectively neutralizes both algebraic and statistical attack vectors.
+Takeaway 5: KV-Cloak completely thwarts all proposed attacks, reducing the quality of reconstructed text to a level statistically indistinguishable from random noise.
+VI-C Inference Accuracy
 TABLE III: Impact of KV-Cloak on inference accuracy (higher is better) across various models, using a block size of 16.
@@
 Qwen2.5-Math-7B 0.620 0.630 0.620 0.632 0.042 0.000
-We tested the impact of KV-Cloak on all experimental models.
-The results in Table III show that across both benchmarks, the impact of KV-Cloak on inference accuracy is negligible.
-This confirms our theoretical design: KV-Cloak is a lossless protection scheme that does not degrade the model’s generation quality or its performance on downstream tasks.
+To rigorously assess model fidelity, we simulate a disaggregated inference service employing a prefill-decode architecture.
+In this pipeline, the KV-cache generated during the prefill phase is secured using either DP or KV-Cloak, transferred, and subsequently utilized by the decode node for token generation.
+We employ two standard benchmarks to measure utility: MMLU [16, 15] for massive multitask knowledge and SQuAD [29] for reading comprehension.
+We evaluated KV-Cloak across all experimental models (results detailed in Table III).
+The empirical data corroborates our theoretical design: KV-Cloak exhibits zero degradation in model performance.
+Unlike Differential Privacy, which forces a trade-off between utility and privacy via noise injection, KV-Cloak leverages reversible linear transformations.
+This ensures that the de-obfuscated attention states are mathematically identical to the plaintext baseline.
+Consequently, the impact on inference accuracy is negligible across both benchmarks, confirming that KV-Cloak is a practically lossless solution.
 Takeaway 6: KV-Cloak is virtually lossless, preserving the model’s fidelity and core utility without any degradation.
-V-C4 Evaluation of Computational Overhead
-Integrating a new cryptographic primitive into a highly-optimized system like vLLM presents a significant engineering challenge.
-Therefore, to accurately assess the performance overhead of KV-Cloak and demonstrate its compatibility with PagedAttention, we employ a dual approach: direct latency measurement combined with architectural analysis.
-To quantify the computational overhead of KV-Cloak, we measured the inference time for the MMLU benchmark under its protection.
-We selected this benchmark specifically to evaluate a worst-case scenario: the MMLU task requires applying KV-Cloak’s obfuscation and de-obfuscation to the entire KV-cache of a long input sequence, yet generates only a single output token.
-This setup maximizes the ratio of cryptographic overhead to inference work.
-Furthermore, our measurements are conservative as they are based on a serial PyTorch-level implementation without any custom CUDA kernel optimizations, and they do not factor in network communication latency.
-TABLE IV: Performance overhead of KV-Cloak (with a block size of 16) on various models during the MMLU benchmark.
-Model Inference Time(s)
-Plaintext KV-Cloak
-LLaMA-7B 10193.1 10649.0 +4.47%
-LLaMA-3.2-1B 1956.8 2155.6 +10.16%
-LLaMA-3.2-3B-Instruct 4525.7 4821.4 +6.53%
-LLaMA-3.1-8B 9797.4 10135.0 +3.45%
-LLaMA-3.1-8B-Distilled 9832.8 10297.7 +4.73%
-Qwen2.5-Math-7B 9184.7 9452.3 +2.91%
-The results of our evaluation across various models are presented in Table IV.
-The findings show that KV-Cloak introduces a modest average latency overhead of approximately 5%.
-Notably, we observe a favorable scaling property: the relative overhead decreases as the model’s hidden size increases.
-In a real-world distributed deployment, the inclusion of network latency between prefill and decode nodes would make the relative impact of KV-Cloak’s computational overhead on the total end-to-end time even lower.
-These empirical results align with our theoretical analysis, confirming that with standard production optimizations, such as operator fusion, the performance impact of KV-Cloak is expected to be minimal.
-Takeaway 7: KV-Cloak introduces negligible computational overhead to the inference pipeline.
-VI Conclusion
+VI-D Evaluation of Performance Overhead
+Figure 9: Micro-benchmark of latency overhead (ms/GB) versus KV-cache size on LLaMA-3.1-8B.
+Computational Overhead.
+To accurately assess the upper bound of performance impact, we employ a conservative, worst-case micro-benchmark simulating a prefill-decode split.
+Measurements are conducted on a serial PyTorch implementation without custom CUDA kernels, isolating the algorithmic overhead.
+As illustrated in Figure 9, latency scales linearly.
+Standard AES encryption incurs a prohibitive cost of 3020.9 ms/GB, nearly rivaling the prefill latency (3438.5 ms/GB) itself.
+In stark contrast, KV-Cloak with operator fusion introduces a negligible overhead of just 15.41 ms/GB.
+This constitutes merely 0.45% of the prefill latency, ranking second only to the insecure DP scheme (4.88 ms/GB).
+In practical deployments, unavoidable network latency would further mask this minimal computation cost, confirming the efficiency of our design.
+Storage Overhead.
+The storage Overhead for KV-Cloak’s secret matrices ($S,M,A$) is negligible.
+The total size, $2n_{\text{kv}}l(b^{2}+\frac{3}{2}d+1)$, is orders of magnitude smaller than the protected cache.
+Quantitatively, for LLaMA-3.1-8B (using block size $b=16$), the overhead is merely 898 KB; even for a 1T-parameter model like Ling-1T [18] (scaling to $b=64$), it remains only 20.9 MB.
+This MB-scale footprint allows all cryptographic keys to securely reside within the limited memory of TEEs with negligible management overhead.
+Takeaway 7: KV-Cloak introduces negligible computational and storage overhead to the inference pipeline.
+VII Discussion and Future Work
+While KV-Cloak provides a solid foundation for protecting the KV-cache, we identify several limitations, which in turn open up exciting avenues for future research.
+Key Management and Hardware Security Integration.
+The security of KV-Cloak fundamentally relies on the confidentiality of its (megabyte-scale) key matrices.
+Currently, we assume these secrets are protected via TEEs.
+Future work should explore the deep integration of KV-Cloak with hardware-level security mechanisms, such as TEEs or confidential GPUs, to construct a more robust defense-in-depth architecture.
+Furthermore, to counter long-term cryptanalysis in persistent services, developing efficient, low-latency key rotation protocols is essential.
+Performance Optimization via Co-Design.
+Although our evaluation shows minimal overhead, hyper-scale deployment demands further optimization.
+At the software level, latency can be masked through the asynchronous generation of One-Time Pad (OTP) matrices, decoupling security operations from the critical inference path.
+At the hardware level, an algorithm-hardware co-design approach—implementing dedicated GPU intrinsics for block-wise permutation and matrix multiplication—could render the obfuscation cost virtually transparent.
+Adaptation to Quantized Models.
+Our current prototype targets floating-point models.
+As the industry pivots toward integer quantization (e.g., INT8 or INT4) for efficiency, extending KV-Cloak is a priority.
+This necessitates designing new, lossless reversible transformations suitable for discrete data types, potentially based on mathematical structures like modular arithmetic.
+VIII Conclusion
 This research exposes a critical security flaw at the heart of modern LLM inference systems: the privacy risk of data leakage from the KV-cache.

### v4

@@
 Crucially, it achieves this robust security with virtually no degradation in model accuracy and minimal performance overhead, offering a practical solution for trustworthy LLM deployment.
+publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2026 23-27 February 2026, San Diego, CA, USA ISBN 979-8-9919276-8-0 https://dx.doi.org/10.14722/ndss.2026.240258 www.ndss-symposium.org
 I Introduction
@@
 However, the storage and potential sharing of the KV-cache introduce significant yet underexplored privacy concerns [41, 39], as illustrated in Figure 1.
-This vulnerability stems from a critical trade-off made in production systems: while end-user communication with the LLM service is typically encrypted (represented by the black lines), the KV-cache itself is almost always processed, transmitted between compute nodes, and persisted in plaintext.
+This vulnerability stems from a critical trade-off made in production systems: while end-user communication with the LLM service is typically encrypted, the KV-cache itself is almost always processed, transmitted between compute nodes, and persisted in plaintext.
 This design choice is a concession to performance, as the unacceptable latency overhead from cryptographically securing the often gigabyte-scale KV-cache would violate the stringent demands of real-time inference.
@@
 This architectural design—not a vulnerability—directly exposes the KV-cache to the Cloud Service Provider (CSP), who inherently possesses it without requiring any additional exploit [43].
-Crucially, the direct correlation between the KV-cache and user inputs [7, 21], combined with the architectural realities that expose it in plaintext, creates a severe and practical attack vector.
+Crucially, the direct correlation between the KV-cache and user inputs [7, 21], combined with its architectural exposure in plaintext, creates a severe and practical attack vector.
 A critical research gap therefore exists in mitigating these risks, necessitating a practical and secure defense mechanism.
@@
 We first explore two relatively direct attack paths.
-The first is the KV-cache Inversion Attack, a method that attempts to leverage known model weight matrices to directly reverse-calculate the input from the KV-cache.
+The first is the KV-cache Inversion Attack, leveraging known model weight matrices to directly reverse-calculate the input from the KV-cache.
 The second, more general-purpose approach is the KV-cache Collision Attack.
@@
 Finally, we introduce a novel, semantic-based KV-cache Injection Attack.
-The core of this attack is to leverage the LLM’s powerful capability to understand and execute instructions.
+This attack leverages the LLM’s powerful capability to understand and execute instructions.
 By appending a specific instruction, such as “Repeat the previous content,” to the end of an intercepted KV-cache context, an adversary can induce the model to “echo” or generalize the core information held within the KV-cache.
@@
 These limitations highlight the urgent need for a novel defense mechanism.
-Therefore, we propose KV-Cloak, a lightweight and secure mechanism for KV-cache obfuscation designed to overcome these challenges.
+Therefore, we propose KV-Cloak, a lightweight and secure KV-cache obfuscation mechanism addressing these challenges.
 At its core, KV-Cloak employs a reversible matrix-based obfuscation scheme that guarantees lossless model accuracy.
@@
 This dynamic permutation prevents adversaries from building stable algebraic relations across multiple queries.
-To further enhance performance, KV-Cloak utilizes operator fusion, an optimization where a portion of the secret obfuscation matrices is algebraically fused into the LLM’s attention layer weights offline.
+To further enhance performance, KV-Cloak utilizes operator fusion, algebraically fusing a portion of the secret obfuscation matrices into the LLM’s attention layer weights offline.
 This shifts the primary computational cost away from the latency-sensitive online inference phase, striking an effective balance between robust security, lossless accuracy, and high performance.
@@
 To mitigate the computational redundancy of autoregressive generation, LLMs utilize the KV-cache mechanism [28].
-Instead of recalculating the Key ($K$) and Value ($V$) matrices for the entire context window at each step, the system persists these intermediate states in memory.
+Instead of recalculating the Key ($K$) and Value ($V$) matrices for the entire context window at each step, the system persists these intermediate states.
 When generating a new token $x_{i}$, the model only computes the current $q_{i},k_{i},v_{i}$, performs attention between $q_{i}$ and the cached history, and subsequently appends the new $(k_{i},v_{i})$ pair to the cache.
@@
 Furthermore, the attack is highly efficient, requiring only a single generation pass.
-This vector underscores that a comprehensive defense must render the KV-cache semantically unintelligible, not just algebraically obfuscated.
+This attack underscores that a comprehensive defense must render the KV-cache semantically unintelligible, not just algebraically obfuscated.
 IV Evaluation of Attacks
@@
 Currently, we assume these secrets are protected via TEEs.
-Future work should explore the deep integration of KV-Cloak with hardware-level security mechanisms, such as TEEs or confidential GPUs, to construct a more robust defense-in-depth architecture.
+Future work should deeply integrate KV-Cloak with hardware-level security mechanisms, such as TEEs or confidential GPUs, to construct a more robust defense-in-depth architecture.
 Furthermore, to counter long-term cryptanalysis in persistent services, developing efficient, low-latency key rotation protocols is essential.
@@
 Our current prototype targets floating-point models.
-As the industry pivots toward integer quantization (e.g., INT8 or INT4) for efficiency, extending KV-Cloak is a priority.
-This necessitates designing new, lossless reversible transformations suitable for discrete data types, potentially based on mathematical structures like modular arithmetic.
+As the industry shifts to integer quantization (e.g., INT8 or INT4) for efficiency, extending KV-Cloak becomes a priority.
+This requires designing new lossless reversible transformations for discrete data types, potentially based on structures like modular arithmetic.
 VIII Conclusion
@@
 It establishes that strong privacy protection can be achieved without sacrificing the utility and efficiency that have made these models so powerful.
+Acknowledgment
+This research is supported in part by the “Pioneer” and “Leading Goose” R&D Program of Zhejiang (2024C01169), and the National Natural Science Foundation of China under Grants (62441238, U2441240).
 Ethic Considerations